A Complete Bibliography of Publications in Reviews of Modern Physics (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.30

Title word cross-reference

1/f [PGFA14]. 2 [SZ19]. 2 + 1 [BTB^{+}10]. 511 [PBB^{+}11]. ^{12} [dGW^{+}17]. ^{16} [dGW^{+}17]. ^{3} [GNSW17]. _{2} [MO11, UBS^{+}16]. _{3} [KKS^{+}12]. _{s} [RP12]. _{s}^{0} [ABL16]. α [THSR17]. α, γ [dGW^{+}17]. B [BMV15]. B_{s}^{0} [ABL19]. t̄t̄ [DG12]. β [HSB^{+}18, VWT15]. PT [KYZ16]. e^{+}e^{−} [DESS11]. η [BM19, BM19]. f(R) [SF10a]. N_{c} [MS15]. p [Ama15]. pp [A^{+}11]. s [KGBA11]. sp^{2} [MSBD16].

-based [MSBD16]. -particle [THSR17]. -symmetric [KYZ16]. -type [Ama15].

125 [BH14].

65 [Mer17].

81 [EHM14, Nam10]. 82 [KC19, Pow10b, TS10b]. 83 [CHT11b, HW11a, KMT11a, SH11b, UHS^{+}17]. 84 [AgCMK12b]. 86 [BCP^{+}14b, TSNK14b, ZJB14b]. 88 [ABL19]. 89 [Giu19].

90 [GHM^{+}22].

Abrikosov [VGG18]. Accelerating [Sch12, Rie12]. acceleration [Per12]. accelerators [CPL^{+}13, DZD^{+}18, HSXZ14, NBD^{+}14]. Achievement [Wit18]. acoustics [FY11]. acoustofluidics [FY11]. across [BRBC^{+}12, FZ12]. active [MJR^{+}13]. adaptive [SF10b]. adherent [SS13]. Adiabatic [AL18, VRSB17]. adiabaticity [GORK^{+}19]. advection [ABB^{+}17]. against [SÁ16]. Alfvén [CZ16]. algebraic [CvD10, DG17]. algorithms [CvD10]. allowed [HSB^{+}18]. Alternatives [CGT16]. amorphous [BB11, ERT19, NFMB18]. ampere [PSM^{+}13]. amplification [CDG^{+}10]. Amplifying [NJBN12]. Analogies [SRHM10]. analogs [LRW^{+}12]. analysis [ES10, EFS15]. analytical [HSB^{+}18]. Anderson [Var15]. angle [MHP^{+}19]. angular [LBTY17]. anisotropies [KAAB^{+}18]. anisotropy [DC17]. annihilation [GYS10, PBB^{+}11, TM13]. Anniversaries [ST18]. Announcement [Ano18]. Annual [FLS13]. Anomalous [KKS19, NSO^{+}10, BM19]. anomaly [BH13]. ansatz [GBL13]. Antiferroelectric [TGC10]. Antiferromagnetic [BMT^{+}18, Dai15]. apertures [GVMMEK10]. Application [KVNF12, Mou19]. applications [ATE^{+}14, CHT11a, CHT11b, CBTW17, DC17, DMM^{+}14, GORK^{+}19, KKS^{+}12, KSLUS18, MMY^{+}12, PRV19, SGOMF13]. approach [MSH^{+}12]. APS [Mey17, Wit18]. Area [ECP10]. article [Wit18]. Artificial [DGJÖ11, NMS13]. aspects [GT10, Hin12]. assembled [SW10]. Astromaterial [CH17]. astronomy [A^{+}13]. astrophysical [HL11]. astrophysics [AD10, KZ11, LFR19]. astrophysics-relevant [LFR19]. Asymmetries [ASAJPV15]. atmospheric [Kaj16]. atom [CGI^{+}19, ES10, SW10, TJG^{+}19]. Atomic [CLW^{+}19, AHV13, BLHE12, CB17, CJC10, DD11, FSRD13, HSP10, HW11a, HW11b, LBY^{+}15, PSO^{+}18, SSdRG11]. atomically [WCG^{+}18]. atomistic [CCSM19]. atoms [CDGT^{+}18, CFRMS19, CDS19, DGJÖ11, HNS13, LMD^{+}13, Met17, RR15, RDBE13, SWM10, SBD^{+}18, TBC^{+}16]. attachment [GYS10]. Attosecond [PNB15]. autoionization [ACM10]. axion [BZ19]. Axions [KC10, KC19].

B [ABL16]. Background [Nak15, KAAB^{+}18]. bands [CDS19]. bang [CFOY16]. barrier [BBS12]. Baryon [KR10, MS15]. baryons [OSZ18]. based [Dai15, DDGS15, DZD^{+}18, MSBD16, RR15, SSdRG11]. Basics [DMM^{+}14]. basis [DC11]. Bayesian [FS13, vT11]. Beam [HSXZ14, Ama15]. Beams [LBTY17, GIF^{+}19]. BEH [Eng14]. behavior [MD18]. Bell [Mer16, Mer17, BCP^{+}14a, BCP^{+}14b, Mer93]. benefits [SESE10]. bent [JLS18]. bent-shaped [JLS18]. Bernal [TS10b, TS10a]. Berry [XCN10]. Beta [AHV13]. Beta-delayed [AHV13]. Bethe [GBL13]. between [HSP10, SRHM10]. Beyond [ISG12, ASG^{+}16, DeG16, GMN^{+}19, LRW^{+}12, OANR^{+}19, RHT^{+}18, TS10a, TS10b, VRSB17]. Big [CFOY16]. billiards [JS17]. binaries [CBKvM10]. binding [DZ10]. biology [BMS19, FKLW19, KT15]. Biophysical [Gol14]. birth [ACM10, Lal14]. bits [SPM^{+}10]. Black [CBKvM10, Bam17, GEG10, Har16, KZ11]. Black-hole [CBKvM10]. blue [Aka15, Nak15]. body [AABS19, GGPR17, HNS13]. bones [BK10]. bootstrap [PRV19]. Born [TBC^{+}16]. Bose [DHY10, SKU13, ZJB14a, ZJB14b]. boson [BH14, Eng14, SRHM10]. bosons [CCG^{+}11]. bound [GYS10]. boundary [Har13, MBC^{+}12]. bounds [dRDTZ17]. box [Har13]. brain [TB18]. breaking [Nam09, Nam10, SH13]. buffer [Ama15]. built [CDGT^{+}18]. bunched [GIF^{+}19]. bundles [PHC10]. burning [CZ16, dGW^{+}17]. bursts [MM12].

C [dGW^{+}17]. calculations [FGH^{+}14]. calculus [Wes14]. caliber [PGLD13]. calorimetry [LLW18, SWK^{+}16]. candidates [Bam17]. carbon [LKS^{+}15, MPG13, MSBD16]. Carlo [CGP^{+}15, GML^{+}11]. Carrier [UHS^{+}11, UHS^{+}17]. case [MO11, Nam09, Nam10]. Casimir [MD18, WDT^{+}16]. cat [Win13]. Cavity [AKM14, RR15, RDBE13]. Cavity-based [RR15]. cavity-generated [RDBE13]. CCD [Boy10, Smi10]. cells [Bet15, SS13]. Center [GEG10]. centuries [Kao10]. chain [A^{+}11]. chains [CLW^{+}19]. change [McD16]. channels [CGLM14]. chaos [CW15, MRW10]. chaotic [ABB^{+}17]. characterization [MBSR^{+}18]. Charge [LPT10, Phi10]. charged [KSLUS18]. chemical [AAL^{+}10]. chemistry [BMS19, VRSB17]. Chief [Mey17]. chiral [BH13]. choice [sMKZ16]. chronoscopy [PNB15]. circuits [NJBN12, XAYN13]. classical [Car14, DS12, Har13, MBC^{+}12, SF10b]. classical-quantum [MBC^{+}12]. Classification [CTSR16]. clathrates [TSNK14a, TSNK14b]. clocks [DK11, KSLUS18, LBY^{+}15]. closed [PSSV11]. Cluster [EDM13, FMBT^{+}10, FW13, GEG10, GGPR17]. clustering [FHKE^{+}18]. clusters [HGH^{+}12, VGG18]. CNO [A^{+}11]. coarse [CCSM19]. CODATA [MTN12]. coexistence [HW11a, HW11b]. coherence [FS13, SAP17]. Coherent [BBF^{+}16, FSRD13]. Cold [RDBE13, TJG^{+}19, VCTL^{+}12, HNS13]. collapse [BLS^{+}13, Bur13]. Collective [MD18, SE11]. Collider [KNS^{+}19]. colliders [ASAJPV15, FGN10]. collisions [DESS11, Shu17]. colloidal [ZMB17]. colloids [MD18]. Colloquium [AABS19, AgCMK12a, AgCMK12b, APV12, AWBF^{+}16, ACM10, BFL^{+}14, BSVH18, BMV15, BP12, BZ19, BLPV16, BK10, BR19, Bur13, BLW14, CHT11a, CH17, CFM^{+}14, CDGT^{+}18, CLW^{+}19, DGJÖ11, DG17, DS12, DK11, DZ10, DMM^{+}14, DM10, DD11, ERT19, ECP10, EF15, FKT15, GMN^{+}19, Gol14, GK18, Gra15, GWY^{+}18, HNS13, HL11, HK10, Hoh10, HGH^{+}12, KKS19, KT15, KVNF12, KKY18, LLP19, LSK14, LRW^{+}12, LL19, MJT15, Met17, Muñ18, MO11, NJBN12, OANR^{+}19, OBN16, Per10, Phi10, PSSV11, RWF17, Rur10, SZ19, SH13, Sch17, SH11a, SE11, SPM^{+}10, SG10, SW10, SKKG10, SAP17, SÁ16, TB18, THSR17, UBS^{+}16, VMU^{+}13, WCG^{+}18, WAC^{+}16, Wen17, Wes14, WG11, ZMB17, ZC13, vHZ10, A^{+}13, AHV13, CHT11b, EDM13, FSRD13, FLS13, LMD^{+}13, NMS13, SH11b]. color [ACC^{+}14]. columns [Gra15]. common [Sca12]. communication [BCMdW10]. Comparison [HL11]. Compass [NvdB15]. Complex [SGOMF13, LB16, PSCVV15, RG17]. complexes [JM18]. Complexity [ST11, BCMdW10, TGC10, Wes14]. Compound [EBD^{+}12]. Compound-nuclear [EBD^{+}12]. compounds [Ste11]. computation [AL18]. Concepts [GORK^{+}19, JWN^{+}14, KT15]. condensate [THSR17]. condensation [DHY10, ZJB14a, ZJB14b]. condensed [BDB^{+}17, CCG^{+}11, CCSM19]. conditions [MMPC12]. conduction [MPG13]. confinement [OBC^{+}14]. conformal [HHSV17, PRV19]. connection [BM19]. considerations [Var15]. consistent [Hoh10]. constants [MTN12]. constraints [LL19]. Continuous [GML^{+}11]. Continuous-time [GML^{+}11]. continuum [RWF17]. Control [LB16, TB18, RG17, Sun14]. Controlling [Har13]. controversial [AWBF^{+}16]. Core [KA12, Bur13]. core-collapse [Bur13]. correction [Ter15]. Corrections [Ano18, SF13]. correlated [BAvdM^{+}11, MSH^{+}12, SKKG10]. correlation [BLHE12]. correlations [CFM^{+}14, HMPW17, MBC^{+}12, RHT^{+}18]. cosmic [KAAB^{+}18, LSS11, Per12]. cosmological [CG10]. cosmology [DS11]. Coulomb [NL16]. counting [EHM09, EHM14]. coupled [Shu17, SGOMF13]. Coupling [SPMS17, FDLR^{+}19]. CP [KC19, ABL16, ABL19, BFJ12, KC10]. CPT [KR11]. critical [MD18]. Criticality [Muñ18]. cross [A^{+}11, EBD^{+}12, FZ12, Nam09, Nam10, TBC^{+}16]. Crystal [MPLS10, MTA^{+}17, TSNK14a, TSNK14b]. Crystalline [ACC^{+}14]. crystals [AgCMK12a, AgCMK12b, CB17, JLS18, TGC10]. cuprates [AFG10]. Current [KUP^{+}12, Kam19, PSM^{+}13]. curved [TVN10]. cycles [A^{+}11].

D [RP12, SZ19]. dark [Bar10, BH18, FLS13]. Data [KR11]. decay [VWT15]. decays [CEN^{+}12, PKGR12, RP12, ZC13]. Defect [TM13]. defects [AgCMK12a, AgCMK12b, FGH^{+}14, Kos17]. definition [PSM^{+}13]. Deformation [NFMB18]. degenerate [SE11]. delay [SGOMF13]. delay-coupled [SGOMF13]. Delayed [sMKZ16, AHV13]. Delayed-choice [sMKZ16]. Delocalized [JM18]. demixing [KAHN10]. Densities [EDM13]. Density [Jon15]. dependent [JWN^{+}14]. deposited [Ama15]. description [HSB^{+}18]. design [HSXZ14]. Designing [NMS13]. detection [Adh14]. detectors [AD10]. developments [BEJR14]. device [JWN^{+}14]. devices [Dor10]. Diagnostics [DZD^{+}18]. diagram [ANT16]. Diagrammatic [RHT^{+}18]. dichalcogenides [WCG^{+}18]. Dielectric [CW15, NBD^{+}14]. diffusion [FSRD13]. Dilute [DO14, SBK^{+}10]. dimension [GBL13]. dimensional [AMV18, CCG^{+}11, GWY^{+}18, ISG12, MSBD16, RWF17, SAHR11, SKKG10]. diodes [Nak15, RTLL13]. dipole [CFRMS19, HvNCR10]. Dirac [AMV18]. directed [EFS15]. Disclination [AgCMK12a, AgCMK12b]. Discord [MBC^{+}12]. Discovery [Kaj16, BH14, Wei18]. Discrete [AF10]. disk [Sel14]. distant [Sch12]. distribution [HR10]. distributions [EDM13]. DNA [Dor10]. Domain [CSRS12]. Domains [JS17]. doped [AFG10, DSF^{+}15, Phi10]. doping [Ama15]. dots [LMS15, UMA^{+}13]. drag [NL16]. drifting [UBS^{+}16]. driven [BLHE12, FMBT^{+}10, FY11, SH13]. droplets [SRHM10]. drum [ST11]. drums [GT10]. Dual [LLW18, RCR12]. Dual-readout [LLW18]. due [MD18]. Dynamic [Sag11]. dynamical [ATE^{+}14, Muñ18, RDBE13, RHT^{+}18, RCR12, SPMS17]. Dynamics [Pow10a, SGOMF13, dVA17, BLPV16, Car14, CG10, Dai15, DeG16, FMBT^{+}10, KVNF12, PSSV11, SKU13, Sun14, TB18, UHS^{+}11, UHS^{+}17, Pow10b].

early [Smi10]. Earth [BK10]. earthquakes [KHK^{+}12]. ecological [ASG^{+}16]. Editor [Mey17]. Editorial [KAB^{+}19, Kam19, Mey17, ST18, Spr14]. Edwards [BMHM18]. EeV [FZ12]. effect [NSO^{+}10]. effects [BSVH18, CGLM14, SVW^{+}15, XCN10]. efficient [Nak15]. Einstein [ZJB14b, DHY10, ZJB14a]. elastic [PHC10]. elastoplastic [NFMB18]. Electric [CFRMS19, BKRB15]. electric-field [BKRB15]. Electrodynamics [BAvdM^{+}11]. electromagnetic [Bam17, GS15]. electromagnetism [LLP19]. Electron [Giu17, KUP^{+}12, LBTY17, Ama15, AFG10, BAvdM^{+}11, DZD^{+}18, GIF^{+}19, PSM^{+}13, PMR16, SE11, SKKG10, TSNK14a, TSNK14b, UBS^{+}16, dA10, Giu19]. electron-crystal [TSNK14a, TSNK14b]. electron-doped [AFG10]. Electron-Electron [KUP^{+}12]. Electronic [CCSM19, Goe11, SAHR11, SW10, LRW^{+}12, RS10, Rur10, XCN10]. electronics [ZDM^{+}13]. electrons [BBB^{+}12, HSXZ14]. electrophoresis [Dor10]. electroweak [HNW12, SF13]. elegantly [Kam19]. Elementary [HvNCR10, AvVD^{+}11, CGT16]. elements [GMN^{+}19, GMKOP12]. Emergent [CFM^{+}14, OANR^{+}19]. emission [GIF^{+}19, PBB^{+}11]. emitting [Nak15, RTLL13]. energetic [CZ16]. energy [Ama15, A^{+}13, FGN10, FZ12, HNW12, LLP19, LSS11]. engineering [KVNF12]. enhanced [BAB^{+}18]. ensembles [HSP10, PSO^{+}18, SSdRG11]. entangled [BMV15]. Entanglement [Nis18, AABS19, BAB^{+}18, ECP10, PCL^{+}12, Win13, Wit18]. Entropic [CBTW17]. entropy [ECP10, Nis18, PGLD13]. environmental [SÁ16]. Epidemic [PSCVV15]. epigenetics [CBC^{+}16]. equation [WAC^{+}16]. equilibrium [MPLS10, SH13]. era [CGI^{+}19]. Erratum [BM16, CHT11b, EHM14, Giu19, GHM^{+}22, KC19, KKR16, Mer16, Mer17, UHS^{+}17, ZJB14b]. error [BLW14, Ter15]. Essential [SPM^{+}10]. Ettore [ACM10]. eV [FZ12]. Evading [Hig14]. events [DG12]. evidence [OSZ18]. evolution [KAHN10, LL19, NS11, Sel14]. evolving [ERT19]. Exceptional [Wit18]. excitations [AvVD^{+}11, FW13, HMPW17, HvNCR10, RS10, dA10]. Exciton [DHY10]. Exciton-polariton [DHY10]. Excitons [WCG^{+}18, JM18]. exoplanets [LL19]. expansion [CG10, Per12, Sch12]. Experiment [TM13].

Experimental [Nag17, OSZ18, SWK^{+}16, BLS^{+}13]. Experiments [Sag11, TSNK14a, TSNK14b, CGI^{+}19, GBL13, sMKZ16]. explored [JWN^{+}14]. Exploring [LFR19, Har13]. extension [Boy10]. Extreme [Mou19, MMPC12]. Extremely [DMHK12].

facilities [LFR19]. Failed [KKS19]. Failure [BK10, PHC10]. Fano [MFK10]. far [SH13]. Fascinated [Aka15]. fast [Kao10]. Femtosecond [CPL^{+}13]. Fermi [GBL13]. Fermion [Wit16, MSH^{+}12, SRHM10]. fermions [Bee15, EF15]. ferromagnetic [DO14]. ferromagnets [BBGK16a, BBGK16b]. fertilization [Nam09, Nam10]. Feshbach [CGJT10]. few [GGPR17]. few-body [GGPR17]. fiber [MTA^{+}17, PHC10]. field [ANT16, ATE^{+}14, BSVH18, BKRB15, Goe11, HHSV17, KKY18, PZ10, RHT^{+}18, RL10, Wit18]. fields [RG17]. fifty [BR19]. filaments [Gra15, Pow10a, Pow10b]. films [KKS^{+}12]. final [NW14]. First [FGH^{+}14, SBK^{+}10, BBF^{+}16, Giu17, Giu19, KT15, MM12]. First-principles [FGH^{+}14, SBK^{+}10]. fission [AHV13]. five [BBF^{+}16]. Flatland [Nov11]. flavor [AF10, McD16]. flavors [BTB^{+}10]. flight [Sun14]. flow [DD11, JBSB^{+}16, LRW^{+}12, NFMB18, SWK^{+}16, Var15]. flows [ST11]. Fluctuation [VGG18, BDG^{+}15, CHT11a, CHT11b, EHM09, EHM14]. fluctuations [EHM09, EHM14]. fluid [Pow10a, Pow10b]. fluids [CC13, SE11, SKKG10]. followed [Ama15]. forces [HNS13, Met17]. formation [GGPR17, KPZ^{+}19, ST11]. Forward [KAB^{+}19]. Foundations [CHT11a, CHT11b, Moe15]. Fractional [LLP19, Wes14]. fracture [KHK^{+}12]. fragility [FSD^{+}18]. free [PMR16]. free-electron [PMR16]. freely [Hel15]. frequency [HL11, LB19]. friction [KHK^{+}12, VMU^{+}13]. frontier [HNW12]. Frontiers [ABB^{+}17]. frustration [NMS13, OANR^{+}19]. function [BLS^{+}13]. Functional [MSH^{+}12, Jon15]. functions [HR10, MMY^{+}12, SPMS17]. fundamental [DMHK12, KSLUS18, MTN12]. fusion [A^{+}11, BEJR14]. future [Jon15, Kao10, MJT15].

Ga [JWN^{+}14]. Galactic [GEG10]. galaxies [Sel14]. Galaxy [CCdJ^{+}14, PBB^{+}11]. gamma [MM12]. GaN [Ama15]. gases [CCG^{+}11, CGJT10, GBL13, MCD^{+}18, SKU13, VCTL^{+}12]. gauge [DGJÖ11]. Gaussian [WPGP^{+}12]. gedanken [sMKZ16]. generated [OBN16, RDBE13]. Generating [Str19]. generators [HCGE17]. Genetic [KAHN10]. genetics [NS11]. Geometry [Gra15]. GeV [BH14]. Ginzburg [RL10]. glass [AWBF^{+}16, BB11, TSNK14a, TSNK14b]. glasses [PZ10]. global [CG10, ES10]. glue [BM19]. gluon [Shu17]. Goldstone [Hig14]. Google [EFS15]. graining [CCSM19]. granular [BMHM18, Gol14]. Graphene [BFL^{+}14, KUP^{+}12, Nov11, Gei11, Goe11, Per10, SAHR11]. Gravitational [Adh14, A^{+}13, Bar18, CBKvM10, Tho18, Wei18]. Graviton [dRDTZ17, GN10]. gravity [BBS12, Hin12, SF10a]. Gripped [DZ10]. group [MSH^{+}12, Nis18]. groups [RCR12]. Growth [Ama15, EDM13].

H [UBS^{+}16]. Hadron [DESS11, HH10, KNS^{+}19, ASAJPV15, FH12]. Hadronic [GHM^{+}18, RP12, NW14, GHM^{+}22]. Hadronization [Alb10]. Hadrons [APV12]. Hall [HHSV17, NSO^{+}10, SVW^{+}15]. hard [PZ10, TS10a, TS10b]. hard-particle [TS10a, TS10b]. harvesting [JM18]. Hearing [GT10]. Heat [DD11, LRW^{+}12]. Heavy [SESE10, BEJR14, CNRS18, OSZ18, Shu17]. Heavy-ion [SESE10, BEJR14]. helium [MMPC12, Var15, dGW^{+}17]. HERA [NW14]. Hermitian [CW15]. heteronanotubes [AAL^{+}10]. Hidden [Car14, Mer93, MO11, Mer16, Mer17]. Hierarchies [HHSV17]. Higgs [BH14, CGT16]. High [GK18, HSB^{+}18, A^{+}13, DMHK12, DMPZ12, FGN10, FKT15, LLP19, MCD^{+}18, RS10, Str19]. high-energy [A^{+}13, FGN10, LLP19]. high-intensity [DMHK12, Str19]. higher [BBS12]. higher-spin [BBS12]. Highly [KSLUS18]. historical [SF13]. History [BH18, Smi10]. hole [Bam17, CBKvM10, GEG10]. holes [Har16, KZ11]. Holography [Nis18]. Homochirality [SH13]. Hoyle [THSR17]. Hybrid [MTA^{+}17, XAYN13, TJG^{+}19]. Hydrodynamics [MJR^{+}13]. hydrogen [ES10, MMPC12]. hypothesis [SH11a, SH11b].

Ice [BRBC^{+}12, OANR^{+}19, NMS13]. icefields [BRBC^{+}12]. identification [TM13]. Identifying [Phi10]. II [A^{+}11, Bar18, RL10]. III [Tho18]. Imaging [BBB^{+}12, Moe15, NMS13]. implementations [FSD^{+}18]. Implications [PGFA14, dGW^{+}17]. improved [BTB^{+}10]. impurity [GML^{+}11]. induced [HTB^{+}17]. Inelastic [RS10, AvVD^{+}11]. inference [vT11]. Influence [CG10]. information [Har16, PGFA14, SWM10, SÁ16, WPGP^{+}12]. infrared [KAAB^{+}18]. InGaN [Nak15]. initial [DESS11]. Insect [Sun14]. Insights [NFMB18, SPMS17]. instabilities [GMKOP12, SW10]. insulators [HK10, Phi10, QZ11]. integrals [Wit16]. intense [OBN16]. intensity [DMHK12, Str19]. interacting [PSSV11, RWF17, XAYN13]. interaction [FDLR^{+}19, Sca12, SPMS17]. Interactions [KUP^{+}12, DG17, DMHK12, FPP^{+}10, GS15, Giu17, Giu19, GMP17, GYS10, MD18, SE11, WDT^{+}16]. Interface [HTB^{+}17, BMS19, HSP10, Sch13]. Interface-induced [HTB^{+}17]. interfaces [CFM^{+}14, DC17, Sag11]. Interfacing [LMS15]. interference [HGH^{+}12]. interferometry [Adh14, PCL^{+}12]. Intermittent [BLMV11]. Interplay [TGC10]. intertwined [FKT15]. intracellular [BN13]. Introduction [CDG^{+}10, Hoh10, Per10]. invention [Nak15, Smi10]. investigation [UMA^{+}13]. Invited [Wit18]. Ion [BKRB15, BEJR14, CNRS18, SESE10, Shu17, TJG^{+}19]. Ion-trap [BKRB15]. Ionic [SZ19]. ionization [BLHE12]. ions [DM10, KSLUS18, SPM^{+}10]. iron [Dai15, Ste11]. iron-based [Dai15]. irradiation [Ama15]. isospectrality [GT10]. itself [BK10].

Jammed [TS10a, BMHM18, TS10b]. jamming [PZ10]. Jeans [VGG18]. Jerusalem [Har16]. Jet [CNRS18, KNS^{+}19]. John [Mer16, Mer17, Mer93]. journal [Lal14]. journeys [Aka15]. junctions [DD11].

Kaon [CEN^{+}12]. kaonic [CGI^{+}19]. Kepler [TS10b, TS10a]. keV [PBB^{+}11]. kinematics [CG10].

laboratory [LFR19]. Landau [DMPZ12, RL10]. large [MS15, KNS^{+}19]. Laser [FMBT^{+}10, HSXZ14, LMD^{+}13, Adh14, BLHE12, CPL^{+}13, DMHK12, NBD^{+}14]. Laser-driven [FMBT^{+}10, BLHE12]. laser-plasma [CPL^{+}13]. lasers [PMR16, SGOMF13]. Lattice [DeG16, GMKOP12, BDY18, DK11, FH12]. lattices [CDGT^{+}18, KMT11a, KMT11b]. laws [ECP10]. layer [Ama15]. learning [CCC^{+}19]. Lecture [Aka15, Ama15, Bar18, Bet15, Eng14, Hal17, Hel15, Hig14, Kos17, Moe15, Nak15, Nam10, Tho18, Wei18, Boy10, Gei11, Har13, Kaj16, Kao10, McD16, Mou19, Nam09, Nov11, Per12, Rie12, Sch12, Smi10, Str19, Win13]. lectures [Har16]. Leptonic [BFJ12]. level [LPT10]. levels [DMPZ12]. Lévy [ZDK15]. LHC [GS18, GMP17]. Life [Sch13, LL19]. lifetime [BH13, WG11]. Light [BBF^{+}16, FH12, GVMMEK10, RG17, Aka15, CC13, CGI^{+}19, DZ10, FDLR^{+}19, FHKE^{+}18, HSP10, Hel15, JM18, LMD^{+}13, Met17, Mou19, Nak15, RTLL13]. light-emitting [RTLL13]. light-harvesting [JM18]. light-matter [FDLR^{+}19]. LIGO [Bar18, Tho18, Wei18]. limits [GN10, PKGR12]. Linac [BBF^{+}16]. linear [KAHN10, SSdRG11]. lines [JS17]. Liquid [AD10, AgCMK12a, AgCMK12b, ISG12, JLS18, TGC10, ZKN17]. liquids [ISG12, MCD^{+}18]. lived [HMPW17]. living [Muñ18, ZMB17]. local [BBB^{+}12, CG10]. localization [AABS19]. localized [DG17, MMY^{+}12]. Long [FPP^{+}10]. Looking [KAB^{+}19, KAAB^{+}18, SG10]. loops [AgCMK12a, AgCMK12b]. Lorentz [KR11]. low [Ama15, MSBD16]. low-dimensional [MSBD16]. low-energy [Ama15]. low-temperature [Ama15]. Luttinger [ISG12].

Machine [CCC^{+}19]. Macroscopic [BDG^{+}15, FSD^{+}18]. Magnetic [FW13, GGG10, HvNCR10, MHP^{+}19, YKJ10, ANT16, DC17, Goe11, KKR10, KKR16, NMS13, RL10, SBK^{+}10]. magnetism [MO11, SKU13, HTB^{+}17]. magnetohydrodynamics [LFR19]. magnets [ZJB14a, ZJB14b]. magnon [ZC13]. Majorana [ACM10, Bee15, EF15]. man [Boy10]. Manipulating [LRW^{+}12]. manipulation [HSXZ14, KKR10, KKR16]. Many [AABS19]. Many-body [AABS19]. Markovian [BLPV16, dVA17]. mass [GN10, MJT15, UBS^{+}16, dRDTZ17]. masses [CCdJ^{+}14, FH12]. massive [GEG10, Hin12, MM12]. Materials [MBSR^{+}18, Nov11, WDT^{+}16, BAvdM^{+}11, BB11, BDB^{+}17, GWY^{+}18, MPG13, SZ19]. Mathematical [GT10]. matrices [MRW10]. matrix [Bee15, EFS15]. matter [BMHM18, BH18, BDB^{+}17, CCG^{+}11, CDGT^{+}18, CTSR16, CCSM19, ERT19, FDLR^{+}19, FLS13, Hal17, KPZ^{+}19, LLP19, MJR^{+}13, Nag17, Sag11, Wen17, ZMB17]. Maximally [MMY^{+}12]. maximum [PGLD13]. Mean [PZ10, ATE^{+}14, BLW14, RHT^{+}18].

Mean-field [PZ10, ATE^{+}14]. measurement [BLW14, CDG^{+}10]. measurements [BH13, BAB^{+}18, BKRB15, CNRS18, EBD^{+}12]. Measures [FSD^{+}18, MBC^{+}12]. Measuring [Per12, WAC^{+}16]. mechanics [ASG^{+}16, BMHM18, EWKK14]. mechanism [Eng14]. mechanisms [SPMS17]. Medal [Wit18]. media [FSRD13, Gol14, RG17]. medium [HH10]. meets [RG17]. membranes [Pow10a, Pow10b]. memories [Ter15]. Memory [KPZ^{+}19, CGLM14]. meson [RP12]. mesons [BM19, BMV15, OSZ18]. mesoscale [VMU^{+}13]. Mesoscopic [BMS19, RG17]. metal [DC17, WCG^{+}18]. metal/oxide [DC17]. Metallic [BBGK16a, GMKOP12, BBGK16b]. metals [DSF^{+}15, KKS19]. metamaterials [LSK14]. metastable [VCTL^{+}12]. methods [CGP^{+}15, GORK^{+}19, GML^{+}11, SF10b]. metrology [PSO^{+}18]. Mg [Ama15]. microcavities [CW15]. microfabricated [Dor10]. Microfluidics [FY11]. microprobes [MBSR^{+}18]. Microscale [FY11]. Microscopic [FHKE^{+}18]. microscopy [Moe15, OBC^{+}14, dA10, vHZ10]. mixing [AF10]. Mn [JWN^{+}14]. Mod [AgCMK12b, ABL19, BBGK16b, BM16, BCP^{+}14b, CHT11b, EHM14, Giu19, GHM^{+}22, HW11a, KMT11a, KC19, KKR16, Mer16, Mer17, Nam10, Pow10b, SH11b, TSNK14b, TS10b, UHS^{+}17, ZJB14b]. Model [CW15, BH14, DG17, DeG16, HNW12, CEN^{+}12]. Modeling [BM14, KVNF12, VMU^{+}13, CCSM19, BM16]. Models [BLS^{+}13, AF10, BN13, GML^{+}11, KGBA11, KAHN10, NFMB18, NvdB15]. Modern [Lal14, CGI^{+}19, HSXZ14, MPLS10]. modes [HvNCR10, KZ11, Phi10]. modulation [FLS13]. molecular [CB17, DD11, Tie13]. molecule [GYS10, Moe15]. molecules [Bet15, BK10, CFRMS19, GHM^{+}18, GHM^{+}22, HGH^{+}12, JLS18, SBD^{+}18, TBC^{+}16]. moments [CFRMS19]. momentum [LBTY17]. Monte [CGP^{+}15, GML^{+}11]. morphologies [EDM13]. motion [HvNCR10]. motivations [NvdB15]. Mott [Phi10]. Multiboson [GMP17]. multicellular [KVNF12]. multifrequency [Met17]. Multimessenger [A^{+}13]. Multiparticle [DS12]. Multiphoton [PCL^{+}12]. multiple [BLHE12]. My [Rie12].

nanobubbles [LZ15]. nanodecomposition [DSF^{+}15]. nanodroplets [LZ15]. nanoelectronics [CSRS12]. Nanoplasmas [OBN16]. nanoprobes [MBSR^{+}18]. nanoscale [FPP^{+}10, MFK10, SZ19, VMU^{+}13]. nanoscopic [CDGT^{+}18]. Nanoscopy [Hel15]. nanostructured [MPG13]. nanostructures [LMS15, MSBD16, OBC^{+}14, Sch17]. nanotubes [LKS^{+}15, MPG13]. nanowires [Rur10]. natural [JM18]. near [BKRB15, KAAB^{+}18]. near-infrared [KAAB^{+}18]. nematic [AgCMK12a, AgCMK12b]. networks [BM14, BM16, DM10, EFS15, PSCVV15, RR15, TB18]. Neutral [ASG^{+}16, BH13, DGJÖ11]. Neutrino [FZ12, GS15, McD16, AF10, Kaj16]. neutrinos [A^{+}13, McD16]. neutron [DS11, LMD^{+}13, MHP^{+}19, VWT15, WAC^{+}16, WG11]. neutron-rich [LMD^{+}13]. Nobel [Aka15, Ama15, Bar18, Bet15, Boy10, Eng14, Gei11, Hal17, Har13, Hel15, Hig14, Kaj16, Kao10, Kos17, McD16, Moe15, Mou19, Nak15, Nam09, Nam10, Nov11, Per12, Rie12, Sch12, Smi10, Str19, Tho18, Wei18, Win13]. noble [VCTL^{+}12]. Nodal [JS17]. noise [BKRB15, CDG^{+}10, PGFA14, SÁ16]. Non [BLPV16, CW15, dVA17]. non-Hermitian [CW15]. Non-Markovian [BLPV16, dVA17]. nonclassical [PSO^{+}18]. Nonequilibrium [ATE^{+}14, BSVH18, DMPZ12, EHM09, FKLW19, PSSV11, EHM14]. Nonlinear [KYZ16, LSK14, SE11, FMBT^{+}10, KMT11a, KMT11b]. nonlocal [RHT^{+}18]. Nonlocality [BCMdW10, BCP^{+}14a, BCP^{+}14b]. Nonperturbative [BTB^{+}10]. Nonstandard [OSZ18]. Normalization [ES10]. Note [AgCMK12b, ABL19, BBGK16b, BCP^{+}14b, HW11a, KMT11a, Nam10, Pow10b, SH11b, TSNK14b, TS10b]. Nuclear [KGBA11, UMA^{+}13, CH17, CGP^{+}15, EF15, EBD^{+}12, GHM16, GEG10, HH10, MRW10, PKGR12, VWT15]. nuclei [AHV13, CJC10, CFRMS19, FHKE^{+}18, HNS13, HvNCR10, HW11a, HW11b, LMD^{+}13]. Nucleon [HMPW17, HR10, ABHM13]. nucleonic [HvNCR10]. nucleosynthesis [CFOY16]. number [HCGE17]. numerical [CBKvM10, PRV19, SPM^{+}10].

O [dGW^{+}17]. Observation [McD16]. observations [KGBA11, Sch12]. Observatory [McD16]. Odd [LB19]. Odd-frequency [LB19]. offshoots [Var15]. Oganesson [GMN^{+}19]. One [CCG^{+}11, ISG12, GBL13, RWF17]. One-dimensional [ISG12, RWF17]. open [BLPV16, dVA17]. Optical [DZ10, KSLUS18, LBY^{+}15, dA10, DK11, KKR10, KKR16, Met17, RR15, RDBE13, SG10, Str19, UMA^{+}13]. Optically [GNSW17]. optics [Bet15, SSdRG11]. optimal [Gra15]. optomechanics [AKM14]. orbital [LBTY17]. order [Dai15, KKR10, KKR16, KT15, MO11]. orders [FKT15]. organic [RTLL13]. origins [Jon15]. oscillations [Kaj16]. other [XAYN13]. oxide [CFM^{+}14, DC17]. oxides [GGG10].

packing [Gra15]. packings [TS10a, TS10b]. pair [ASAJPV15]. pairing [RCR12, Sca12]. paradigm [ISG12]. particle [AD10, DS11, EF15, Nam09, Nam10, OANR^{+}19, Sch13, SWK^{+}16, THSR17, TS10a, TS10b]. particles [CZ16, CFRMS19, ZMB17]. partonic [LPT10]. partons [Alb10]. passage [VRSB17]. passing [GVMMEK10]. past [Kao10]. pasta [CH17]. path [Rie12, Wit16]. pattern [ST11]. patterns [BRBC^{+}12]. periodic [KKY18]. Perpendicular [DC17]. Persistent [Sch17]. perspective [BB11, RTLL13, SF13, WDT^{+}16]. Perspectives [Bur13, KUP^{+}12, MJT15, AFG10]. perturbations [ES10]. Phase [ANT16, CJC10, Kos17, XCN10]. phases [Wen17, Wit16]. phenomena [BBB^{+}12, DMPZ12, JWN^{+}14, KKY18, MPG13, SZ19, HTB^{+}17]. Phonon [TSNK14a, Giu17, Giu19, TSNK14b]. Phonon-glass [TSNK14a, TSNK14b]. Phononic [GWY^{+}18]. Phononics [LRW^{+}12]. photocontrol [Moe15]. photoelectron [BLHE12]. photoemission [PNB15]. Photon [GN10]. photonic [LMS15, MTA^{+}17]. photonic-crystal [MTA^{+}17]. photonics [OPA^{+}19, SGOMF13]. photons [BBB^{+}12, CDGT^{+}18, Har13, LMS15, RR15, RWF17]. Phys [AgCMK12b, ABL19, BBGK16b, BM16, BCP^{+}14b, CHT11b, EHM14, Giu19, GHM^{+}22, HW11a, KMT11a, KC19, KKR16, Mer16, Nam10, Pow10b, SH11b, TSNK14b, TS10b, UHS^{+}17, ZJB14b]. Phys. [Mer17]. Physical [DC11, LL19, MSBD16, SESE10, AAL^{+}10, BBB^{+}12, CCC^{+}19, GT10, KKS^{+}12, MTN12, NvdB15, Ano18, ST18]. Physics [CZ16, DK11, DO14, JLS18, SH11a, SH11b, SS13, AD10, BMS19, BZ19, CW15, Car14, CGP^{+}15, CNRS18, CBC^{+}16, DG12, DS11, ERT19, EF15, FKLW19, GHM16, GS15, GGPR17, HHSV17, KGBA11, KHK^{+}12, KA12, KT15, KSLUS18, LLP19, MRW10, Mou19, Nam09, Nam10, PMR16, PGLD13, RCR12, SBD^{+}18, Sch13, SF13, UMA^{+}13, VRSB17, vT11, Lal14]. physics-chemistry-biology [BMS19]. pion [BH13, HR10]. plane [CFM^{+}14, TBC^{+}16]. plane-wave [TBC^{+}16]. plants [JBSB^{+}16]. Plasma [DDGS15, CPL^{+}13, DZD^{+}18, Shu17]. plasma-based [DZD^{+}18]. plasmas [CZ16, SE11]. plasmons [DG17]. point [AgCMK12a, AgCMK12b, FGH^{+}14]. polariton [DHY10]. polaritons [FSRD13]. polarized [GNSW17, OBC^{+}14]. polymer [BM14, BM16]. pores [SZ19]. portraits [JS17]. Positron [GYS10, PBB^{+}11, TM13]. Positron-molecule [GYS10]. positrons [DDGS15]. potentials [DGJÖ11, RDBE13]. power [LFR19]. precision [HSB^{+}18, SF13]. Predictive [CCSM19]. Present [CFOY16]. pressure [GK18, MCD^{+}18, RS10]. Principles [PGLD13, FGH^{+}14, Giu17, Giu19, Gol14, LB16, SBK^{+}10]. prism [SG10]. probes [BBB^{+}12]. probing [LMD^{+}13]. problem [KC10, KC19]. problems [CvD10]. process [KGBA11]. processes [BRBC^{+}12, BDY18, PSCVV15, PHC10]. production [ASAJPV15, DESS11, GS18]. Progress [AFG10]. prominence [Jon15]. propagating [Hel15, Phi10]. properties [AAL^{+}10, CFM^{+}14, Goe11, GWY^{+}18, HH10, KKS^{+}12, MMPC12, MSBD16, Per10, Rur10, Sag11, Wit18, XCN10]. propulsion [Gol14]. Protecting [SÁ16]. protection [DC11]. Proteins [ERT19]. proton [UBS^{+}16]. PRX [Spr14]. Publications [Ano18]. Publisher [AgCMK12b, ABL19, BBGK16b, BCP^{+}14b, HW11a, KMT11a, Nam10, Pow10b, SH11b, TSNK14b, TS10b]. pulsed [LFR19]. pulsed-power [LFR19]. pulses [Str19]. puzzle [BR19]. pyrochlore [GGG10].

QCD [ANT16, BTB^{+}10, BH13, BDY18, FH12, MS15]. quantitative [NS11]. Quantum [BAB^{+}18, BLW14, CHT11a, CHT11b, CGP^{+}15, CGLM14, CC13, CJC10, CDGT^{+}18, CvD10, CG19, DRC17, DM10, FS13, GAN14, HSP10, HHSV17, HCGE17, HGH^{+}12, LKS^{+}15, PSO^{+}18, SWM10, SSdRG11, SAP17, Ter15, ZKN17, AL18, BMV15, BBGK16a, BBGK16b, BLPV16, Car14, CB17, CTSR16, CDG^{+}10, DS12, DMHK12, DMM^{+}14, EHM09, EHM14, FSD^{+}18, GML^{+}11, Hal17, Har16, Har13, Hoh10, ISG12, JS17, LLP19, LMS15, MBC^{+}12, NJBN12, OBC^{+}14, PGFA14, PSSV11, RR15, SRHM10, SE11, SPM^{+}10, SKU13, SÁ16, UMA^{+}13, WPGP^{+}12, Wen17, Wit18, XAYN13, ZJB14a, ZJB14b, ZDM^{+}13, dVA17]. Quantum-Bayesian [FS13]. Quantum-enhanced [BAB^{+}18]. quantum-entangled [BMV15]. quantum-to-classical [DS12]. quantum-topological [Wen17]. quark [ASAJPV15, DG12, GS18, Shu17]. quark-gluon [Shu17]. quarks [BTB^{+}10, HMPW17]. Quasinormal [KZ11].

radiation [Adh14, Bam17, DESS11, DC11, KAAB^{+}18, OBN16]. Radiative [SF13]. Radioactive [PKGR12]. radiobiological [SESE10]. raising [Win13]. Raman [VRSB17]. Random [Bee15, Gei11, KT15, MRW10, HCGE17]. Random-matrix [Bee15]. range [FPP^{+}10]. ratio [UBS^{+}16]. ray [AvVD^{+}11, MM12, MBSR^{+}18, PMR16, RS10, WAC^{+}16]. Rayleigh [VGG18]. rays [CPL^{+}13, LSS11]. reaction [EBD^{+}12, dGW^{+}17]. reactions [BEJR14, MRW10]. readout [LLW18]. Real [EWKK14]. Real-space [EWKK14]. realization [Ama15]. realizations [sMKZ16]. recommended [MTN12]. reconnection [YKJ10]. reconstruction [SF10b]. refined [PSM^{+}13]. regimes [FDLR^{+}19]. region [HR10]. related [MPG13, MBC^{+}12]. relations [BLW14, CHT11a, CHT11b, CBTW17]. relativity [CBKvM10]. relevant [LFR19]. renormalization [EWKK14, MSH^{+}12, Nis18]. repeaters [SSdRG11]. research [Kam19, KA12, Wit18]. resolution [Bet15, Moe15]. resolved [UHS^{+}11, UHS^{+}17, vHZ10]. resonance [BR19]. resonances [BDY18, CGJT10, MS15, MFK10]. Resonant [AvVD^{+}11, GYS10]. resource [CG19, SAP17]. Rev [AgCMK12b, ABL19, BBGK16b, BM16, BCP^{+}14b, CHT11b, EHM14, Giu19, GHM^{+}22, HW11a, KMT11a, KC19, KKR16, Mer16, Nam10, Pow10b, SH11b, TSNK14b, TS10b, UHS^{+}17, ZJB14b]. Rev. [Mer17]. reversal [BMV15]. Review [ST18, Ano18, ST18]. Reviews [Lal14]. rich [LMD^{+}13]. Riemann [SH11b, SH11a]. rise [Jon15]. RMP [KAB^{+}19]. road [GK18]. Role [Spr14, DS11]. room [GK18]. root [BLW14]. root-mean-square [BLW14]. Roper [BR19]. Rotating [MM12, ST11]. rotating-drum [ST11]. routes [RHT^{+}18]. rule [OANR^{+}19]. Rydberg [SWM10].

Sand [Kao10]. Sap [JBSB^{+}16]. sapphire [Ama15]. scalar [Eng14]. Scaled [TBC^{+}16]. scales [FZ12]. scaling [Muñ18]. Scanning [OBC^{+}14, vHZ10]. Scattering [BDY18, AvVD^{+}11, MHP^{+}19, RG17, RS10]. scene [Lal14]. Schrödinger’s [Win13]. science [CH17, DDGS15, FPP^{+}10, Nag17]. sciences [CCC^{+}19]. scientific [Lal14]. Search [SBD^{+}18, UBS^{+}16, BLMV11]. Searches [FGN10, BH14]. sections [A^{+}11, EBD^{+}12, FZ12, TBC^{+}16]. Secular [Sel14]. segregation [ST11]. self [Gol14, SW10]. self-assembled [SW10]. self-propulsion [Gol14]. semiconductor [Sch17]. semiconductors [DO14, DSF^{+}15, SBK^{+}10, SGOMF13, TM13, UHS^{+}11, UHS^{+}17]. semiflexible [BM14, BM16]. semimetals [AMV18]. Send [Kao10]. sensing [DRC17]. Shape [APV12, HW11b, HSB^{+}18, HW11a]. shaped [JLS18]. shapes [CJC10, GT10]. short [HMPW17]. short-lived [HMPW17]. Shortcuts [GORK^{+}19]. Si [MO11]. Silicon [ZDM^{+}13, Rur10]. simplicity [TGC10]. Simulation [CB17, GAN14]. simulations [BTB^{+}10, BMS19]. Single [Bet15, GS18, Moe15, PSM^{+}13, LMS15, RR15]. Single-electron [PSM^{+}13]. Single-molecule [Moe15]. size [EDM13]. small [MHP^{+}19]. small-angle [MHP^{+}19]. soft [BDB^{+}17, MJR^{+}13, Nag17, Sag11]. soft-matter [Nag17]. Solar [A^{+}11, McD16]. solid [EF15, PGFA14]. solid-state [EF15, PGFA14]. Solids [MCD^{+}18, AMV18, FGH^{+}14, NFMB18]. soliton [SG10]. Solitons [KMT11a, KMT11b]. solution [BR19]. Some [Var15]. Source [BBF^{+}16]. sources [PSM^{+}13]. space [DC11, EWKK14]. spectrometry [MJT15]. spectroscopy [BFL^{+}14, KR10, Moe15, UHS^{+}11, UHS^{+}17, VGG18]. spectrum [HSB^{+}18]. sphere [PZ10]. Spin [JWN^{+}14, OBC^{+}14, SVW^{+}15, ABHM13, BBS12, BSVH18, CLW^{+}19, Dai15, NMS13, Sch17, UMA^{+}13, ZKN17]. Spin-dependent [JWN^{+}14]. Spin-polarized [OBC^{+}14]. spin-splitting [BSVH18]. spin-two [BBS12]. Spinodal [DSF^{+}15]. Spinor [SKU13]. spintronic [DO14]. spintronics [BMT^{+}18]. splitting [BSVH18]. Spontaneous [Nam09, Nam10, ZC13]. square [BLW14]. SrRuO [KKS^{+}12]. Stability [Sun14, PKGR12]. staggered [BTB^{+}10]. Standard [BH14, DeG16, HNW12, CEN^{+}12]. standards [HL11]. star [GEG10, WAC^{+}16]. stars [BZ19, MM12]. state [DESS11, EF15, KA12, NW14, PGFA14, THSR17, WAC^{+}16]. states [FSD^{+}18, GYS10, PSO^{+}18, ZKN17]. Statistical [ASG^{+}16, KHK^{+}12, NS11, BMHM18, EWKK14, PGLD13]. statistics [EHM09, EHM14, JS17]. Status [KUP^{+}12, RTLL13, THSR17, CFOY16]. Steady [KA12]. Steady-state [KA12]. stellar [KGBA11, dGW^{+}17]. stepping [KAHN10]. Stimulated [VRSB17, GIF^{+}19]. stimulated-superradiant [GIF^{+}19]. Stimulating [NJBN12]. Stochastic [BN13]. stone [KAHN10]. story [Nak15]. Strangeness [GHM16]. strategies [BLMV11]. stress [BDB^{+}17]. string [KZ11, Sch13]. Strong [CFM^{+}14, KKY18, Met17, Goe11, KC10, KC19]. Strong-field [KKY18]. Strongly [RWF17, Shu17, SKKG10]. Structural [Rur10]. Structure [KKS^{+}12, ABHM13, THSR17]. structures [BRBC^{+}12, DO14, MFK10]. studied [UHS^{+}11, UHS^{+}17]. studies [AvVD^{+}11]. substructure [KNS^{+}19]. subwavelength [GVMMEK10]. succinctly [Kam19]. Sudbury [McD16]. sugar [JBSB^{+}16]. super [Bet15, Moe15]. super-resolution [Bet15, Moe15]. Superconducting [XAYN13, NJBN12]. Superconductivity [Ste11, GK18, LB19, MO11]. superconductors [ACC^{+}14, Bee15, BSVH18, Dai15, FKT15, KKS19, QZ11, RL10, Sca12]. supercontinuum [SG10]. superfluid [Var15]. Superheavy [GMN^{+}19]. supernova [Bur13]. supernovae [Per12, Sch12]. Superposition [Win13]. superpositions [DS12]. Superradiant [GIF^{+}19]. Supersolids [BP12]. supersymmetry [FGN10]. Surface [LZ15, DG17]. surfaces [BKRB15, CLW^{+}19, EDM13, MPLS10, TVN10]. surpasses [BBS12]. surrogate [EBD^{+}12]. symmetric [KYZ16]. Symmetries [SKU13, AF10, Car14, CTSR16]. Symmetry [SH13, VWT15, LPT10, Nam09, Nam10, RCR12]. synchrotron [MBSR^{+}18]. Synthesizing [Kam19]. system [ABL16, ABL19]. systems [ASG^{+}16, BLPV16, CW15, CCG^{+}11, DMHK12, EHM09, EHM14, KYZ16, KVNF12, KKY18, LB16, MSH^{+}12, Muñ18, PSSV11, SRHM10, SH13, TJG^{+}19, XAYN13, dVA17].

tables [KR11]. Takes [Spr14, Lal14]. Tale [ST18]. techniques [DDGS15, HHSV17, PRV19]. temperature [Ama15, FKT15, GK18]. terahertz [UHS^{+}11, UHS^{+}17]. terrestrial [HL11]. Testing [Bam17]. Tests [HNW12, BLS^{+}13, DeG16, SWK^{+}16]. Tevatron [DG12, GS18]. textures [Sch17]. their [CBTW17, DG17, sMKZ16]. theorem [Hig14]. theorems [EHM09, EHM14, Mer93, Mer16, Mer17]. Theoretical [BB11, Hin12]. Theories [BLHE12, BLS^{+}13, CG19, SF10a]. Theory [FKT15, MMY^{+}12, NvdB15, PRV19, ATE^{+}14, ASG^{+}16, Bee15, BDG^{+}15, Bur13, HHSV17, Hoh10, Jon15, KT15, KZ11, PZ10, RHT^{+}18, RL10, Sag11, SBK^{+}10, Sch13, TSNK14a, TSNK14b, TM13, Wit18]. therapy [SESE10]. Thermal [MPG13, GWY^{+}18]. thermalization [AABS19]. thermoelectric [TSNK14a, TSNK14b]. thermoelectricity [DD11]. thin [KKS^{+}12, WCG^{+}18]. thread [Sca12]. Three [HNS13, AMV18]. Three-body [HNS13]. three-dimensional [AMV18]. Time [BMV15, vHZ10, GML^{+}11, UHS^{+}11, UHS^{+}17]. Time-resolved [vHZ10, UHS^{+}11, UHS^{+}17]. Time-reversal [BMV15]. timing [WAC^{+}16]. tissue [KVNF12]. tokamak [KA12]. tools [SPM^{+}10]. Top [DG12, ASAJPV15, GS18]. top-quark [GS18]. Topological [CDS19, Hal17, HK10, Kos17, OPA^{+}19, QZ11, Bee15, CTSR16, Wen17, Wit16]. Track [SF10b]. traits [NS11]. transition [BB11, DS12, DC17, DSF^{+}15, KT15, WCG^{+}18]. transitions [AWBF^{+}16, CJC10, Kos17]. Transport [SKKG10, BN13, JBSB^{+}16, LKS^{+}15, Per10, Rur10, SAHR11]. trap [BKRB15, DDGS15]. trap-based [DDGS15]. Trapped [SPM^{+}10, DM10, VCTL^{+}12]. travel [DC11]. trends [MJT15]. tumor [SESE10]. tunneling [OBC^{+}14, vHZ10]. tutorial [Wes14]. twisted [Gra15]. Two [ST18, BBS12, GWY^{+}18, Mer93, Mer16, Mer17, SAHR11, SKKG10]. two-dimensional [GWY^{+}18, SAHR11]. type [Ama15, RL10].

ultracold [CCG^{+}11, CGJT10, CDS19]. Ultrafast [KKR10, KKR16]. Ultrahigh [LSS11]. ultrashort [Str19]. ultrasonics [FY11]. Ultrastrong [FDLR^{+}19]. uncertainty [BLW14, CBTW17, NJBN12]. unconventional [Sca12]. underlying [BLS^{+}13]. Understanding [DMM^{+}14, CB17]. undulatory [Gol14]. Universal [GGPR17, SPMS17]. Universe [Sch12, Tie13, Bar10, Rie12]. unsolved [MO11]. URu [MO11]. using [DG12, Per12, WAC^{+}16].

vacuum [NJBN12]. valence [HR10]. values [DMM^{+}14, MTN12]. variables [Mer93, Mer16, Mer17]. vertex [SF10b]. via [Ama15, DESS11, FY11]. view [BRBC^{+}12, Boy10, CFM^{+}14, MPLS10, Wes14]. violation [ABL16, ABL19, BMV15, BFJ12, KR11]. violations [VWT15]. viscous [Pow10a, Pow10b]. voices [Kao10]. vortex [VGG18]. Vortices [SRHM10, TVN10, LBTY17].

Waals [WDT^{+}16]. walk [Gei11]. walks [ZDK15]. wall [CSRS12]. Wannier [MMY^{+}12]. Water [AWBF^{+}16]. wave [BLS^{+}13, CW15, RG17, TBC^{+}16]. wave-function [BLS^{+}13]. waves [A^{+}13, Bar18, CBKvM10, CZ16, KYZ16, Tho18, VGG18, Wei18]. weak [DMM^{+}14]. Weyl [AMV18]. where [BP12]. White [RTLL13]. window [GS15]. wires [SW10]. within [HMPW17]. without [BAB^{+}18].

X [AvVD^{+}11, CPL^{+}13, MBSR^{+}18, PMR16, RS10, WAC^{+}16]. X-ray [AvVD^{+}11, MBSR^{+}18, PMR16, RS10, WAC^{+}16]. xenon [AD10].

year [BR19]. Years [ST18, BBF^{+}16, MJT15]. Yield [BDB^{+}17].

Zoo [Wen17].

References

Adelberger:2011:SFC

[A^{+}11] E. G. Adelberger et al. Solar fusion cross sections. II. The pp chain and CNO cycles. Reviews of Modern Physics, 83(1): 195–245, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

             aps.org/doi/10.1103/RevModPhys.83.195;   http://rmp.
             aps.org/abstract/RMP/v83/i1/p195_1.

Ando:2013:MAG

[A^{+}13] Shin’ichiro Ando et al. Colloquium: Multimessenger astronomy with gravitational waves and high-energy neutrinos. Reviews of Modern Physics, 85(4):1401–1420, October 2013. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

             0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
             85.1401;   http://rmp.aps.org/abstract/RMP/v85/i4/p1401_

Abanin:2019:CMB

[AABS19] Dmitry A. Abanin, Ehud Altman, Immanuel Bloch, and Maksym Serbyn. Colloquium: Many-body localization, thermalization, and entanglement. Reviews of Modern Physics, 91 (2):021001–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
021001.

Ayala:2010:PCP

[AAL^{+}10] Paola Ayala, Raul Arenal, Annick Loiseau, Angel Rubio, and Thomas Pichler. The physical and chemical properties of heteronanotubes. Reviews of Modern Physics, 82(2): 1843–1885, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1843;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1843_1.

Aref:2017:FCA

[ABB^{+}17] Hassan Aref, John R. Blake, Marko Budisić, Silvana S. S. Cardoso, Julyan H. E. Cartwright, Herman J. H. Clercx, Kamal El Omari, Ulrike Feudel, Ramin Golestanian, Emmanuelle Gouillart, GertJan F. van Heijst, Tatyana S. Krasnopolskaya, Yves Le Guer, Robert S. MacKay, Vyacheslav V. Meleshko, Guy Metcalfe, Igor Mezić, Alessandro P. S. de Moura, Oreste Piro, Michel F. M. Speetjens, Rob Sturman, Jean-Luc Thiffeault, and Idan Tuval. Frontiers of chaotic advection. Reviews of Modern Physics, 89(2):025007–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.025007.

Aidala:2013:SSN

[ABHM13] Christine A. Aidala, Steven D. Bass, Delia Hasch, and Gerhard K. Mallot. The spin structure of the nucleon. Reviews of Modern Physics, 85(2):655–691, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.85.655;
http://rmp.aps.org/abstract/RMP/v85/i2/p655_1.

Artuso:2016:CVB

[ABL16] Marina Artuso, Guennadi Borissov, and Alexander Lenz. CP violation in the B_{s}^{0} system. Reviews of Modern Physics, 88 (4):045002–??, October/December 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://www.w3.org/1998/Math/MathML. See publisher’s note

[ABL19] correcting equations 47–49..

Artuso:2019:PNC

[ABL19] Marina Artuso, Guennadi Borissov, and Alexander Lenz. Publisher’s note: CP violation in the B_{s}^{0} system [Rev. Mod. Phys. 88, 45002 (2016)]. Reviews of Modern Physics, 91(4):049901– ??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.w3.org/

1998/Math/MathML. See [ABL16].

Anglani:2014:CCS

[ACC^{+}14] Roberto Anglani, Roberto Casalbuoni, Marco Ciminale, Nicola Ippolito, Raoul Gatto, Massimo Mannarelli, and Marco Ruggieri. Crystalline color superconductors. Reviews of Modern Physics, 86(2):509–??, April/June 2014. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.86.509.

Arimondo:2010:CEM

[ACM10] E. Arimondo, Charles W. Clark, and W. C. Martin. Colloquium: Ettore Majorana and the birth of autoionization. Reviews of Modern Physics, 82(3):1947–1958, July 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.1947;   http://rmp.aps.org/abstract/RMP/v82/i3/p1947_
1.

Aprile:2010:LXD

[AD10] E. Aprile and T. Doke. Liquid xenon detectors for particle physics and astrophysics. Reviews of Modern Physics, 82 (3):2053–2097, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2053;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2053_1.

Adhikari:2014:GRD

[Adh14] Rana X. Adhikari. Gravitational radiation detection with laser interferometry. Reviews of Modern Physics, 86(1):121–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.121.

Altarelli:2010:DFS

[AF10] Guido Altarelli and Ferruccio Feruglio. Discrete flavor symmetries and models of neutrino mixing. Reviews of Modern Physics, 82(3):2701–2729, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2701;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2701_1.

Armitage:2010:PPE

[AFG10] N. P. Armitage, P. Fournier, and R. L. Greene. Progress and perspectives on electron-doped cuprates. Reviews of Modern Physics, 82(3):2421–2487, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2421;   http://rmp.aps.org/abstract/RMP/v82/i3/p2421_
1.

Alexander:2012:CDL

[AgCMK12a] Gareth P. Alexander, Bryan Gin ge Chen, Elisabetta A. Matsumoto, and Randall D. Kamien. Colloquium: Disclination loops, point defects, and all that in nematic liquid crystals. Reviews of Modern Physics, 84(2):497–514, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.84.497;   http://rmp.aps.
org/abstract/RMP/v84/i2/p497_1. See publisher’s note

[AgCMK12b].

Alexander:2012:PNC

[AgCMK12b] Gareth P. Alexander, Bryan Gin ge Chen, Elisabetta A. Matsumoto, and Randall D. Kamien. Publisher’s note: Colloquium: Disclination loops, point defects, and all that in nematic liquid crystals [Rev. Mod. Phys. 84, 497 (2012)]. Reviews of Modern Physics, 84(3):1229, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1229;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1229_1. See [AgCMK12a].

Andreyev:2013:BDF

[AHV13] Andrei N. Andreyev, Mark Huyse, and Piet Van Duppen. Colloquium: Beta-delayed fission of atomic nuclei. Reviews of Modern Physics, 85(4):1541–1559, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.85.
1541;   http://rmp.aps.org/abstract/RMP/v85/i4/p1541_
1.

Akasaki:2015:NLF

[Aka15] Isamu Akasaki. Nobel lecture: Fascinated journeys into blue light. Reviews of Modern Physics, 87(4):1119–??, October/ December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.1119.

Aspelmeyer:2014:CO

[AKM14] Markus Aspelmeyer, Tobias J. Kippenberg, and Florian Marquardt. Cavity optomechanics. Reviews of Modern Physics, 86 (4):1391–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.86.
1391.

Albash:2018:AQC

[AL18] Tameem Albash and Daniel A. Lidar. Adiabatic quantum computation. Reviews of Modern Physics, 90(1):015002–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.015002.

Albino:2010:HP

[Alb10] S. Albino. Hadronization of partons. Reviews of Modern Physics, 82(3):2489–2556, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2489;   http://rmp.aps.org/abstract/RMP/v82/i3/p2489_
1.

Amano:2015:NLG

[Ama15] Hiroshi Amano. Nobel lecture: Growth of GaN on sapphire via low-temperature deposited buffer layer and realization of p-type GaN by Mg doping followed by low-energy electron beam irradiation. Reviews of Modern Physics, 87(4):1133–??, October/ December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.1133.

Armitage:2018:WDS

[AMV18] N. P. Armitage, E. J. Mele, and Ashvin Vishwanath. Weyl and Dirac semimetals in three-dimensional solids. Reviews of Modern Physics, 90(1):015001–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.015001.

Anonymous:2018:ACP

[Ano18] Anonymous. Announcement: Corrections in Physical Review publications. Reviews of Modern Physics, 90(1):010002–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.90.010002.

Andersen:2016:PDQ

[ANT16] Jens O. Andersen, William R. Naylor, and Anders Tranberg. Phase diagram of QCD in a magnetic field. Reviews of Modern Physics, 88(2):025001–??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.025001.

Alexandrou:2012:CSH

[APV12] Constantia Alexandrou, Costas N. Papanicolas, and Marc Vanderhaeghen. Colloquium: The shape of hadrons. Reviews of Modern Physics, 84(3):1231–1251, July 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.1231;   http://rmp.aps.org/abstract/RMP/v84/i3/p1231_
1.

Aguilar-Saavedra:2015:ATQ

[ASAJPV15] J. A. Aguilar-Saavedra, D. Amidei, A. Juste, and M. Pérez-Victoria. Asymmetries in top quark pair production at hadron colliders. Reviews of Modern Physics, 87(2):421–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.87.421.

Azaele:2016:SME

[ASG^{+}16] Sandro Azaele, Samir Suweis, Jacopo Grilli, Igor Volkov, Jayanth R. Banavar, and Amos Maritan. Statistical mechanics of ecological systems: Neutral theory and beyond. Reviews of Modern Physics, 88(3):035003–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.035003.

Aoki:2014:NDM

[ATE^{+}14] Hideo Aoki, Naoto Tsuji, Martin Eckstein, Marcus Kollar, Takashi Oka, and Philipp Werner. Nonequilibrium dynamical mean-field theory and its applications. Reviews of Modern Physics, 86(2):779–??, April/June 2014. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.86.779.

Ament:2011:RIX

[AvVD^{+}11] Luuk J. P. Ament, Michel van Veenendaal, Thomas P. Devereaux, John P. Hill, and Jeroen van den Brink. Resonant inelastic X-ray scattering studies of elementary excitations. Reviews of Modern Physics, 83(2):705–767, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.705;
http://rmp.aps.org/abstract/RMP/v83/i2/p705_1.

Amann-Winkel:2016:CWC

[AWBF^{+}16] Katrin Amann-Winkel, Roland Böhmer, Franz Fujara, Catalin Gainaru, Burkhard Geil, and Thomas Loerting. Colloquium: Water’s controversial glass transitions. Reviews of Modern Physics, 88(1):011002–??, January 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.011002.

Braun:2018:QEM

[BAB^{+}18] Daniel Braun, Gerardo Adesso, Fabio Benatti, Roberto Floreanini, Ugo Marzolino, Morgan W. Mitchell, and Stefano Pirandola. Quantum-enhanced measurements without entanglement. Reviews of Modern Physics, 90(3):035006–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.035006.

Bambi:2017:TBH

[Bam17] Cosimo Bambi. Testing black hole candidates with electromagnetic radiation. Reviews of Modern Physics, 89(2):025001–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.89.025001.

Bartelmann:2010:DU

[Bar10] Matthias Bartelmann. The dark Universe. Reviews of Modern Physics, 82(1):331–382, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.331;
http://rmp.aps.org/abstract/RMP/v82/i1/p331_1.

Barish:2018:NLL

[Bar18] Barry C. Barish. Nobel lecture: LIGO and gravitational waves II. Reviews of Modern Physics, 90(4):040502–??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL https://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.040502.

Basov:2011:ECE

[BAvdM^{+}11] D. N. Basov, Richard D. Averitt, Dirk van der Marel, Martin Dressel, and Kristjan Haule. Electrodynamics of correlated electron materials. Reviews of Modern Physics, 83(2): 471–541, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.83.471;   http://rmp.
aps.org/abstract/RMP/v83/i2/p471_1.

Berthier:2011:TPG

[BB11] Ludovic Berthier and Giulio Biroli. Theoretical perspective on the glass transition and amorphous materials. Reviews of Modern Physics, 83(2):587–645, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.587;
http://rmp.aps.org/abstract/RMP/v83/i2/p587_1.

Bonnell:2012:IPP

[BBB^{+}12] Dawn A. Bonnell, D. N. Basov, Matthias Bode, Ulrike Diebold, Sergei V. Kalinin, Vidya Madhavan, Lukas Novotny, Miquel Salmeron, Udo D. Schwarz, and Paul S. Weiss. Imaging physical phenomena with local probes: From electrons to photons. Reviews of Modern Physics, 84(3):1343– 1381, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1343;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1343_1.

Bostedt:2016:LCL

[BBF^{+}16] Christoph Bostedt, Sébastien Boutet, David M. Fritz, Zhirong Huang, Hae Ja Lee, Henrik T. Lemke, Aymeric Robert, William F. Schlotter, Joshua J. Turner, and Garth J. Williams. Linac Coherent Light Source: The first five years. Reviews of Modern Physics, 88(1):015007–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.015007.

Brando:2016:MQF

[BBGK16a] M. Brando, D. Belitz, F. M. Grosche, and T. R. Kirkpatrick. Metallic quantum ferromagnets. Reviews of Modern Physics, 88(2):025006–??, February 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.025006.

Brando:2016:PNM

[BBGK16b] M. Brando, D. Belitz, F. M. Grosche, and T. R. Kirkpatrick. Publisher’s note: Metallic quantum ferromagnets [Rev. Mod. Phys. 88, 25006 (2016)]. Reviews of Modern Physics, 88(3):039901–??, July/September 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.039901.

Bekaert:2012:HHS

[BBS12] Xavier Bekaert, Nicolas Boulanger, and Per A. Sundell. How higher-spin gravity surpasses the spin-two barrier. Reviews of Modern Physics, 84(3):987–1009, July 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.987;   http://rmp.aps.org/abstract/RMP/v84/i3/p987_
1.

Buhrman:2010:NCC

[BCMdW10] Harry Buhrman, Richard Cleve, Serge Massar, and Ronald de Wolf. Nonlocality and communication complexity. Reviews of Modern Physics, 82(1):665–698, January 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.665;   http://rmp.aps.org/abstract/RMP/v82/i1/p665_
1.

Brunner:2014:BN

[BCP^{+}14a] Nicolas Brunner, Daniel Cavalcanti, Stefano Pironio, Valerio Scarani, and Stephanie Wehner. Bell nonlocality. Reviews of Modern Physics, 86(2):419–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.419. See note [BCP^{+}14b].

Brunner:2014:PNB

[BCP^{+}14b] Nicolas Brunner, Daniel Cavalcanti, Stefano Pironio, Valerio Scarani, and Stephanie Wehner. Publisher’s note: Bell nonlocality [Rev. Mod. Phys. 86, 419 (2014)]. Reviews of Modern Physics, 86(2):839–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.839. See [BCP^{+}14a].

Bonn:2017:YSM

[BDB^{+}17] Daniel Bonn, Morton M. Denn, Ludovic Berthier, Thibaut Divoux, and Sébastien Manneville. Yield stress materials in soft condensed matter. Reviews of Modern Physics, 89(3):035005–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.035005.

Bertini:2015:MFT

[BDG^{+}15] Lorenzo Bertini, Alberto De Sole, Davide Gabrielli, Giovanni Jona-Lasinio, and Claudio Landim. Macroscopic fluctuation theory. Reviews of Modern Physics, 87(2):593–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.87.593.

Briceno:2018:SPR

[BDY18] Raúl A. Briceño, Jozef J. Dudek, and Ross D. Young. Scattering processes and resonances from lattice QCD. Reviews of Modern Physics, 90(2):025001–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.025001.

Beenakker:2015:RMT

[Bee15] C. W. J. Beenakker. Random-matrix theory of Majorana fermions and topological superconductors. Reviews of Modern Physics, 87(3):1037–??, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1037.

Back:2014:RDH

[BEJR14] B. B. Back, H. Esbensen, C. L. Jiang, and K. E. Rehm. Recent developments in heavy-ion fusion reactions. Reviews of Modern Physics, 86(1):317–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.317.

Betzig:2015:NLS

[Bet15] Eric Betzig. Nobel lecture: Single molecules, cells, and superresolution optics. Reviews of Modern Physics, 87(4):1153–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.87.
1153.

Branco:2012:LCV

[BFJ12] G. C. Branco, R. González Felipe, and F. R. Joaquim. Leptonic CP violation. Reviews of Modern Physics, 84(2):515– 565, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.84.515;   http://rmp.aps.
org/abstract/RMP/v84/i2/p515_1.

Basov:2014:CGS

[BFL^{+}14] D. N. Basov, M. M. Fogler, A. Lanzara, Feng Wang, and Yuanbo Zhang. Colloquium: Graphene spectroscopy. Reviews of Modern Physics, 86(3):959–??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.959.

Bernstein:2013:NPL

[BH13] A. M. Bernstein and Barry R. Holstein. Neutral pion lifetime measurements and the QCD chiral anomaly. Reviews of Modern Physics, 85(1):49–77, January 2013. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
85.49;   http://rmp.aps.org/abstract/RMP/v85/i1/p49_
1.

Bernardi:2014:SMH

[BH14] Gregorio Bernardi and Matthew Herndon. Standard model Higgs boson searches through the 125 GeV boson discovery. Reviews of Modern Physics, 86(2):479–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.479.

Bertone:2018:HDM

[BH18] Gianfranco Bertone and Dan Hooper. History of dark matter. Reviews of Modern Physics, 90(4):045002:1–045002:32, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.045002.

Buehler:2010:CFM

[BK10] Markus J. Buehler and Sinan Keten. Colloquium: Failure of molecules, bones, and the Earth itself. Reviews of Modern Physics, 82(2):1459–1487, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1459;   http://rmp.aps.org/abstract/RMP/v82/i2/p1459_

Brownnutt:2015:ITM

[BKRB15] M. Brownnutt, M. Kumph, P. Rabl, and R. Blatt. Ion-trap measurements of electric-field noise near surfaces. Reviews of Modern Physics, 87(4):1419–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1419.

Becker:2012:TPC

[BLHE12] Wilhelm Becker, XiaoJun Liu, Phay Jo Ho, and Joseph H. Eberly. Theories of photoelectron correlation in laser-driven multiple atomic ionization. Reviews of Modern Physics, 84 (3):1011–1043, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1011;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1011_1.

Benichou:2011:ISS

[BLMV11] O. Bénichou, C. Loverdo, M. Moreau, and R. Voituriez. Intermittent search strategies. Reviews of Modern Physics, 83 (1):81–129, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.83.81;   http:/
/rmp.aps.org/abstract/RMP/v83/i1/p81_1.

Breuer:2016:CNM

[BLPV16] Heinz-Peter Breuer, Elsi-Mari Laine, Jyrki Piilo, and Bassano Vacchini. Colloquium: Non-Markovian dynamics in open quantum systems. Reviews of Modern Physics, 88(2):021002–??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.88.021002.

Bassi:2013:MWF

[BLS^{+}13] Angelo Bassi, Kinjalk Lochan, Seema Satin, Tejinder P. Singh, and Hendrik Ulbricht. Models of wave-function collapse, underlying theories, and experimental tests. Reviews of Modern Physics, 85(2):471–527, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.85.471;
http://rmp.aps.org/abstract/RMP/v85/i2/p471_1.

Busch:2014:CQR

[BLW14] Paul Busch, Pekka Lahti, and Reinhard F. Werner. Colloquium: Quantum root-mean-square error and measurement uncertainty relations. Reviews of Modern Physics, 86(4):1261–??, April/ June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.86.1261.

Broedersz:2014:MSP

[BM14] C. P. Broedersz and F. C. MacKintosh. Modeling semiflexible polymer networks. Reviews of Modern Physics, 86(3):995– ??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.86.

995. See erratum [BM16].

Broedersz:2016:EMS

[BM16] C. P. Broedersz and F. C. MacKintosh. Erratum: Modeling semiflexible polymer networks [Rev. Mod. Phys. 86, 995 (2014)]. Reviews of Modern Physics, 88(3):039903–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.039903. See [BM14].

Bass:2019:MCA

[BM19] Steven D. Bass and Pawel Moskal. η′ and η mesons with connection to anomalous glue. Reviews of Modern Physics, 91(1): 015003–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.

w3.org/1998/Math/MathML.

Baule:2018:ESM

[BMHM18] Adrian Baule, Flaviano Morone, Hans J. Herrmann, and Hernán A. Makse. Edwards statistical mechanics for jammed granular matter. Reviews of Modern Physics, 90(1):015006–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.015006.

Bernaschi:2019:MSP

[BMS19] Massimo Bernaschi, Simone Melchionna, and Sauro Succi. Mesoscopic simulations at the physics-chemistry-biology interface. Reviews of Modern Physics, 91(2):025004–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.025004.

Baltz:2018:AS

[BMT^{+}18] V. Baltz, A. Manchon, M. Tsoi, T. Moriyama, T. Ono, and Y. Tserkovnyak. Antiferromagnetic spintronics. Reviews of Modern Physics, 90(1):015005–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.015005.

Bernabeu:2015:CTR

[BMV15] J. Bernabéu and F. Martínez-Vidal. Colloquium: Time-reversal violation with quantum-entangled B mesons. Reviews of Modern Physics, 87(1):165–??, January/March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.165.

Bressloff:2013:SMI

[BN13] Paul C. Bressloff and Jay M. Newby. Stochastic models of intracellular transport. Reviews of Modern Physics, 85(1):135–196, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.85.135;   http://rmp.aps.
org/abstract/RMP/v85/i1/p135_1.

Boyle:2010:NLC

[Boy10] Willard S. Boyle. Nobel Lecture: CCD — an extension of man’s view. Reviews of Modern Physics, 82(3):2305– 2306, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2305;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2305_1.

Boninsegni:2012:CSW

[BP12] Massimo Boninsegni and Nikolay V. Prokof’ev. Colloquium: Supersolids: What and where are they? Reviews of Modern Physics, 84(2):759–776, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.759;
http://rmp.aps.org/abstract/RMP/v84/i2/p759_1.

Burkert:2019:CRR

[BR19] Volker D. Burkert and Craig D. Roberts. Colloquium: Roper resonance: Toward a solution to the fifty year puzzle. Reviews of Modern Physics, 91(1):011003–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.011003.

Bartels-Rausch:2012:ISP

[BRBC^{+}12] Thorsten Bartels-Rausch, Vance Bergeron, Julyan H. E. Cartwright, Rafael Escribano, John L. Finney, Hinrich Grothe, Pedro J. Gutiérrez, Jari Haapala, Werner F. Kuhs, Jan B. C. Pettersson, Stephen D. Price, C. Ignacio Sainz-Díaz, Debbie J. Stokes, Giovanni Strazzulla, Erik S. Thomson, Hauke Trinks, and Nevin Uras-Aytemiz. Ice structures, patterns, and processes: a view across the icefields. Reviews of Modern Physics, 84(2):885–944, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.84.885;   http://rmp.
aps.org/abstract/RMP/v84/i2/p885_1.

Bergeret:2018:CNE

[BSVH18] F. Sebastian Bergeret, Mikhail Silaev, Pauli Virtanen, and Tero T. Heikkilä. Colloquium: Nonequilibrium effects in superconductors with a spin-splitting field. Reviews of Modern Physics, 90(4):041001–??, April 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.041001.

Bazavov:2010:NQS

[BTB^{+}10] A. Bazavov, D. Toussaint, C. Bernard, J. Laiho, C. DeTar, L. Levkova, M. B. Oktay, Steven Gottlieb, U. M. Heller, J. E. Hetrick, P. B. Mackenzie, R. Sugar, and R. S. Van de Water. Nonperturbative QCD simulations with 2 + 1 flavors of improved staggered quarks. Reviews of Modern Physics, 82 (2):1349–1417, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1349;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1349_1.

Burrows:2013:CPC

[Bur13] Adam Burrows. Colloquium: Perspectives on core-collapse supernova theory. Reviews of Modern Physics, 85(1):245–261, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.85.245;   http://rmp.aps.
org/abstract/RMP/v85/i1/p245_1.

Braaten:2019:CPA

[BZ19] Eric Braaten and Hong Zhang. Colloquium: The physics of axion stars. Reviews of Modern Physics, 91(4):041002–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.041002.

Cariglia:2014:HSD

[Car14] Marco Cariglia. Hidden symmetries of dynamics in classical and quantum physics. Reviews of Modern Physics, 86(4):1283–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.1283.

Cazorla:2017:SUA

[CB17] Claudio Cazorla and Jordi Boronat. Simulation and understanding of atomic and molecular quantum crystals. Reviews of Modern Physics, 89(3):035003–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.035003.

Cortini:2016:PE

[CBC^{+}16] Ruggero Cortini, Maria Barbi, Bertrand R. Caré, Christophe Lavelle, Annick Lesne, Julien Mozziconacci, and Jean-Marc Victor. The physics of epigenetics. Reviews of Modern Physics, 88(2):025002–??, February 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.025002.

Centrella:2010:BHB

[CBKvM10] Joan Centrella, John G. Baker, Bernard J. Kelly, and James R. van Meter. Black-hole binaries, gravitational waves, and numerical relativity. Reviews of Modern Physics, 82(4):3069– 3119, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.3069;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p3069_1.

Coles:2017:EUR

[CBTW17] Patrick J. Coles, Mario Berta, Marco Tomamichel, and Stephanie Wehner. Entropic uncertainty relations and their applications. Reviews of Modern Physics, 89(1):015002–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.015002.

Carusotto:2013:QFL

[CC13] Iacopo Carusotto and Cristiano Ciuti. Quantum fluids of light. Reviews of Modern Physics, 85(1):299–366, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.299;   http://rmp.aps.org/
abstract/RMP/v85/i1/p299_1.

Carleo:2019:MLP

[CCC^{+}19] Giuseppe Carleo, Ignacio Cirac, Kyle Cranmer, Laurent Daudet, Maria Schuld, Naftali Tishby, Leslie Vogt-Maranto, and Lenka Zdeborová. Machine learning and the physical sciences. Reviews of Modern Physics, 91(4):045002–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.045002.

Courteau:2014:GM

[CCdJ^{+}14] Stéphane Courteau, Michele Cappellari, Roelof S. de Jong, Aaron A. Dutton, Eric Emsellem, Henk Hoekstra, L. V. E. Koopmans, Gary A. Mamon, Claudia Maraston, Tommaso Treu, and Lawrence M. Widrow. Galaxy masses. Reviews of Modern Physics, 86(1):47–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.47.

Cazalilla:2011:ODB

[CCG^{+}11] M. A. Cazalilla, R. Citro, T. Giamarchi, E. Orignac, and M. Rigol. One dimensional bosons: From condensed matter systems to ultracold gases. Reviews of Modern Physics, 83(4): 1405–1466, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1405;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1405_1.

Cipcigan:2019:ECG

[CCSM19] F. S. Cipcigan, J. Crain, V. P. Sokhan, and G. J. Martyna. Electronic coarse graining: Predictive atomistic modeling of condensed matter. Reviews of Modern Physics, 91(2):025003–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.91.025003.

Clerk:2010:IQN

[CDG^{+}10] A. A. Clerk, M. H. Devoret, S. M. Girvin, Florian Marquardt, and R. J. Schoelkopf. Introduction to quantum noise, measurement, and amplification. Reviews of Modern Physics, 82 (2):1155–1208, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1155;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1155_1.

Chang:2018:CQM

[CDGT^{+}18] D. E. Chang, J. S. Douglas, A. González-Tudela, C.-L. Hung, and H. J. Kimble. Colloquium: Quantum matter built from nanoscopic lattices of atoms and photons. Reviews of Modern Physics, 90(3):031002–??, March 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.031002.

Cooper:2019:TBU

[CDS19] N. R. Cooper, J. Dalibard, and I. B. Spielman. Topological bands for ultracold atoms. Reviews of Modern Physics, 91(1): 015005–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
015005.

Cirigliano:2012:KDS

[CEN^{+}12] Vincenzo Cirigliano, Gerhard Ecker, Helmut Neufeld, Antonio Pich, and Jorge Portolés. Kaon decays in the Standard Model. Reviews of Modern Physics, 84(1):399–447, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.84.399;   http://rmp.aps.org/
abstract/RMP/v84/i1/p399_1.

Chakhalian:2014:CEP

[CFM^{+}14] Jak Chakhalian, John W. Freeland, Andrew J. Millis, Christos Panagopoulos, and James M. Rondinelli. Colloquium: Emergent properties in plane view: Strong correlations at oxide interfaces. Reviews of Modern Physics, 86(4):1189–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.1189.

Cyburt:2016:BBN

[CFOY16] Richard H. Cyburt, Brian D. Fields, Keith A. Olive, and Tsung-Han Yeh. Big bang nucleosynthesis: Present status. Reviews of Modern Physics, 88(1):015004–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.015004.

Chupp:2019:EDM

[CFRMS19] T. E. Chupp, P. Fierlinger, M. J. Ramsey-Musolf, and J. T. Singh. Electric dipole moments of atoms, molecules, nuclei, and particles. Reviews of Modern Physics, 91(1):015001–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.015001.

Carrera:2010:IGC

[CG10] Matteo Carrera and Domenico Giulini. Influence of global cosmological expansion on local dynamics and kinematics. Reviews of Modern Physics, 82(1):169–208, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.169;   http://rmp.aps.org/
abstract/RMP/v82/i1/p169_1.

Chitambar:2019:QRT

[CG19] Eric Chitambar and Gilad Gour. Quantum resource theories. Reviews of Modern Physics, 91(2):025001–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.025001.

Curceanu:2019:MEL

[CGI^{+}19] Catalina Curceanu, Carlo Guaraldo, Mihail Iliescu, Michael Cargnelli, Ryugo Hayano, Johann Marton, Johann Zmeskal, Tomoichi Ishiwatari, Masa Iwasaki, Shinji Okada, Diana Laura Sirghi, and Hideyuki Tatsuno. The modern era of light kaonic atom experiments. Reviews of Modern Physics, 91(2):025006– ??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.91.025006.

Chin:2010:FRU

[CGJT10] Cheng Chin, Rudolf Grimm, Paul Julienne, and Eite Tiesinga. Feshbach resonances in ultracold gases. Reviews of Modern Physics, 82(2):1225–1286, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1225;   http://rmp.aps.org/abstract/RMP/v82/i2/p1225_
1.

Caruso:2014:QCM

[CGLM14] Filippo Caruso, Vittorio Giovannetti, Cosmo Lupo, and Stefano Mancini. Quantum channels and memory effects. Reviews of Modern Physics, 86(4):1203–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.1203.

Carlson:2015:QMC

[CGP^{+}15] J. Carlson, S. Gandolfi, F. Pederiva, Steven C. Pieper, R. Schiavilla, K. E. Schmidt, and R. B. Wiringa. Quantum Monte Carlo methods for nuclear physics. Reviews of Modern Physics, 87(3):1067–??, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1067.

Csaki:2016:AEH

[CGT16] Csaba Csáki, Christophe Grojean, and John Terning. Alternatives to an elementary Higgs. Reviews of Modern Physics, 88(4):045001–??, October/December 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.045001.

Caplan:2017:CAS

[CH17] M. E. Caplan and C. J. Horowitz. Colloquium: Astromaterial science and nuclear pasta. Reviews of Modern Physics, 89(4):041002–??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.041002.

Campisi:2011:CQF

[CHT11a] Michele Campisi, Peter Hänggi, and Peter Talkner. Colloquium: Quantum fluctuation relations: Foundations and applications. Reviews of Modern Physics, 83(3):771–791, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.771;   http://rmp.aps.org/
abstract/RMP/v83/i3/p771_1. See erratum [CHT11b].

Campisi:2011:ECQ

[CHT11b] Michele Campisi, Peter Hänggi, and Peter Talkner. Erratum: Colloquium: Quantum fluctuation relations: Foundations and applications [Rev. Mod. Phys. 83, 771 (2011)]. Reviews of Modern Physics, 83(4):1653, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.83.
1653;   http://rmp.aps.org/abstract/RMP/v83/i4/p1653_

1. See [CHT11a].

Cejnar:2010:QPT

[CJC10] Pavel Cejnar, Jan Jolie, and Richard F. Casten. Quantum phase transitions in the shapes of atomic nuclei. Reviews of Modern Physics, 82(3):2155–2212, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2155;   http://rmp.aps.org/abstract/RMP/v82/i3/p2155_

Choi:2019:CAS

[CLW^{+}19] Deung-Jang Choi, Nicolas Lorente, Jens Wiebe, Kirsten von Bergmann, Alexander F. Otte, and Andreas J. Heinrich. Colloquium: Atomic spin chains on surfaces. Reviews of Modern Physics, 91(4):041001–??, April 2019. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.91.041001.

Connors:2018:JMH

[CNRS18] Megan Connors, Christine Nattrass, Rosi Reed, and Sevil Salur. Jet measurements in heavy ion physics. Reviews of Modern Physics, 90(2):025005–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.025005.

Corde:2013:FXR

[CPL^{+}13] S. Corde, K. Ta Phuoc, G. Lambert, R. Fitour, V. Malka, A. Rousse, A. Beck, and E. Lefebvre. Femtosecond X rays from laser-plasma accelerators. Reviews of Modern Physics, 85(1):1–48, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.85.1;
 http://rmp.aps.org/abstract/RMP/v85/i1/p1_1.

Catalan:2012:DWN

[CSRS12] G. Catalan, J. Seidel, R. Ramesh, and J. F. Scott. Domain wall nanoelectronics. Reviews of Modern Physics, 84(1):119–156, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.84.119;   http://rmp.aps.
org/abstract/RMP/v84/i1/p119_1.

Chiu:2016:CTQ

[CTSR16] Ching-Kai Chiu, Jeffrey C. Y. Teo, Andreas P. Schnyder, and Shinsei Ryu. Classification of topological quantum matter with symmetries. Reviews of Modern Physics, 88(3):035005–??, July/ September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.88.035005.

Childs:2010:QAA

[CvD10] Andrew M. Childs and Wim van Dam. Quantum algorithms for algebraic problems. Reviews of Modern Physics, 82(1):1–52, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.82.1;   http://
rmp.aps.org/abstract/RMP/v82/i1/p1_1.

Cao:2015:DMM

[CW15] Hui Cao and Jan Wiersig. Dielectric microcavities: Model systems for wave chaos and non-Hermitian physics. Reviews of Modern Physics, 87(1):61–??, January/March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.61.

Chen:2016:PAW

[CZ16] Liu Chen and Fulvio Zonca. Physics of Alfvén waves and energetic particles in burning plasmas. Reviews of Modern Physics, 88(1):015008–??, January 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.015008.

deAbajo:2010:OEE

[dA10] F. J. García de Abajo. Optical excitations in electron microscopy. Reviews of Modern Physics, 82(1):209–275, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.209;   http://rmp.aps.org/
abstract/RMP/v82/i1/p209_1.

Dai:2015:AOS

[Dai15] Pengcheng Dai. Antiferromagnetic order and spin dynamics in iron-based superconductors. Reviews of Modern Physics, 87(3): 855–??, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.87.
855.

Durante:2011:PBR

[DC11] Marco Durante and Francis A. Cucinotta. Physical basis of radiation protection in space travel. Reviews of Modern Physics, 83 (4):1245–1281, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1245;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1245_1.

Dieny:2017:PMA

[DC17] B. Dieny and M. Chshiev. Perpendicular magnetic anisotropy at transition metal/oxide interfaces and applications. Reviews of Modern Physics, 89(2):025008–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.025008.

Dubi:2011:CHF

[DD11] Yonatan Dubi and Massimiliano Di Ventra. Colloquium: Heat flow and thermoelectricity in atomic and molecular junctions. Reviews of Modern Physics, 83(1):131–155, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.131;   http://rmp.aps.org/
abstract/RMP/v83/i1/p131_1.

Danielson:2015:PTB

[DDGS15] J. R. Danielson, D. H. E. Dubin, R. G. Greaves, and C. M. Surko. Plasma and trap-based techniques for science with positrons. Reviews of Modern Physics, 87(1):247–??, January/ March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.87.247.

DeGrand:2016:LTB

[DeG16] Thomas DeGrand. Lattice tests of beyond standard model dynamics. Reviews of Modern Physics, 88(1):015001–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.015001.

Druzhinin:2011:HPC

[DESS11] V. P. Druzhinin, S. I. Eidelman, S. I. Serednyakov, and E. P. Solodov. Hadron production via e^{+}e^{−} collisions with initial state radiation. Reviews of Modern Physics, 83(4):1545– 1588, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1545;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1545_1.

Deliot:2012:TQP

[DG12] Frédéric Déliot and Douglas A. Glenzinski. Top quark physics at the Tevatron using t̄t̄events. Reviews of Modern Physics, 84(1): 211–252, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.84.211;   http://rmp.
aps.org/abstract/RMP/v84/i1/p211_1.

Davis:2017:CAM

[DG17] T. J. Davis and D. E. Gómez. Colloquium: An algebraic model of localized surface plasmons and their interactions. Reviews of Modern Physics, 89(1):011003–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.011003.

Dalibard:2011:CAG

[DGJÖ11] Jean Dalibard, Fabrice Gerbier, Gediminas Juzeliunas, and Patrik Öhberg. Colloquium: Artificial gauge potentials for neutral atoms. Reviews of Modern Physics, 83(4):1523– 1543, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1523;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1523_1.

deBoer:2017:COR

[dGW^{+}17] R. J. deBoer, J. Görres, M. Wiescher, R. E. Azuma, A. Best, C. R. Brune, C. E. Fields, S. Jones, M. Pignatari, D. Sayre, K. Smith, F. X. Timmes, and E. Uberseder. The C^{12} (α, γ)O^{16} reaction and its implications for stellar helium burning. Reviews of Modern Physics, 89(3):035007–??, March 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://www.w3.org/1998/Math/MathML.

Deng:2010:EPB

[DHY10] Hui Deng, Hartmut Haug, and Yoshihisa Yamamoto. Excitonpolariton Bose–Einstein condensation. Reviews of Modern Physics, 82(2):1489–1537, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1489;   http://rmp.aps.org/abstract/RMP/v82/i2/p1489_
1.

Derevianko:2011:CPO

[DK11] Andrei Derevianko and Hidetoshi Katori. Colloquium: Physics of optical lattice clocks. Reviews of Modern Physics, 83(2):331– 347, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.83.331;   http://rmp.aps.
org/abstract/RMP/v83/i2/p331_1.

Duan:2010:CQN

[DM10] L.-M. Duan and C. Monroe. Colloquium: Quantum networks with trapped ions. Reviews of Modern Physics, 82 (2):1209–1224, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1209;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1209_1.

DiPiazza:2012:EHI

[DMHK12] A. Di Piazza, C. Müller, K. Z. Hatsagortsyan, and C. H. Keitel. Extremely high-intensity laser interactions with fundamental quantum systems. Reviews of Modern Physics, 84 (3):1177–1228, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1177;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1177_1.

Dressel:2014:CUQ

[DMM^{+}14] Justin Dressel, Mehul Malik, Filippo M. Miatto, Andrew N. Jordan, and Robert W. Boyd. Colloquium: Understanding quantum weak values: Basics and applications. Reviews of Modern Physics, 86(1):307–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.307.

Dmitriev:2012:NPH

[DMPZ12] I. A. Dmitriev, A. D. Mirlin, D. G. Polyakov, and M. A. Zudov. Nonequilibrium phenomena in high Landau levels. Reviews of Modern Physics, 84(4):1709–1763, October 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.1709;   http://rmp.aps.org/abstract/RMP/v84/i4/p1709_
1.

Dietl:2014:DFS

[DO14] Tomasz Dietl and Hideo Ohno. Dilute ferromagnetic semiconductors: Physics and spintronic structures. Reviews of Modern Physics, 86(1):187–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.187.

Dorfman:2010:DEM

[Dor10] Kevin D. Dorfman. DNA electrophoresis in microfabricated devices. Reviews of Modern Physics, 82(4):2903– 2947, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2903;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p2903_1.

Degen:2017:QS

[DRC17] C. L. Degen, F. Reinhard, and P. Cappellaro. Quantum sensing. Reviews of Modern Physics, 89(3):035002–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.035002.

deRham:2017:GMB

[dRDTZ17] Claudia de Rham, J. Tate Deskins, Andrew J. Tolley, and Shuang-Yong Zhou. Graviton mass bounds. Reviews of Modern Physics, 89(2):025004–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.025004.

Dubbers:2011:NRC

[DS11] Dirk Dubbers and Michael G. Schmidt. The neutron and its role in cosmology and particle physics. Reviews of Modern Physics, 83(4):1111–1171, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1111;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1111_1.

DeMartini:2012:CMQ

[DS12] Francesco De Martini and Fabio Sciarrino. Colloquium: Multiparticle quantum superpositions and the quantum-to-classical transition. Reviews of Modern Physics, 84(4):1765– 1789, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1765;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1765_1.

Dietl:2015:SNS

[DSF^{+}15] T. Dietl, K. Sato, T. Fukushima, A. Bonanni, M. Jamet, A. Barski, S. Kuroda, M. Tanaka, Pham Nam Hai, and H. Katayama-Yoshida. Spinodal nanodecomposition in semiconductors doped with transition metals. Reviews of Modern Physics, 87(4):1311–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1311.

deVega:2017:DNM

[dVA17] Inés de Vega and Daniel Alonso. Dynamics of non-Markovian open quantum systems. Reviews of Modern Physics, 89(1): 015001–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.89.
015001.

Dholakia:2010:CGL

[DZ10] Kishan Dholakia and Pavel Zemánek. Colloquium: Gripped by light: Optical binding. Reviews of Modern Physics, 82 (2):1767–1791, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1767;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1767_1.

Downer:2018:DPB

[DZD^{+}18] M. C. Downer, R. Zgadzaj, A. Debus, U. Schramm, and M. C. Kaluza. Diagnostics for plasma-based electron accelerators. Reviews of Modern Physics, 90(3):035002–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.035002.

Escher:2012:CNR

[EBD^{+}12] Jutta E. Escher, Jason T. Burke, Frank S. Dietrich, Nicholas D. Scielzo, Ian J. Thompson, and Walid Younes. Compoundnuclear reaction cross sections from surrogate measurements. Reviews of Modern Physics, 84(1):353–397, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.84.353;   http://rmp.aps.org/
abstract/RMP/v84/i1/p353_1.

Eisert:2010:CAL

[ECP10] J. Eisert, M. Cramer, and M. B. Plenio. Colloquium: Area laws for the entanglement entropy. Reviews of Modern Physics, 82 (1):277–306, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.82.277;   http:
//rmp.aps.org/abstract/RMP/v82/i1/p277_1.

Einax:2013:CGS

[EDM13] Mario Einax, Wolfgang Dieterich, and Philipp Maass. Colloquium: Cluster growth on surfaces: Densities, size distributions, and morphologies. Reviews of Modern Physics, 85 (3):921–939, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.85.921;   http://rmp.
aps.org/abstract/RMP/v85/i3/p921_1.

Elliott:2015:CMF

[EF15] Steven R. Elliott and Marcel Franz. Colloquium: Majorana fermions in nuclear, particle, and solid-state physics. Reviews of Modern Physics, 87(1):137–??, January/March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.137.

Ermann:2015:GMA

[EFS15] Leonardo Ermann, Klaus M. Frahm, and Dima L. Shepelyansky. Google matrix analysis of directed networks. Reviews of Modern Physics, 87(4):1261–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1261.

Esposito:2009:NFF

[EHM09] Massimiliano Esposito, Upendra Harbola, and Shaul Mukamel. Nonequilibrium fluctuations, fluctuation theorems, and counting statistics in quantum systems. Reviews of Modern Physics, 81(4):1665–1702, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1665;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1665_1. See erratum

[EHM14].

Esposito:2014:ENF

[EHM14] Massimiliano Esposito, Upendra Harbola, and Shaul Mukamel. Erratum: Nonequilibrium fluctuations, fluctuation theorems, and counting statistics in quantum systems [Rev. Mod. Phys. 81, 1665 (2009)]. Reviews of Modern Physics, 86(3):1125– ??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.86.

1125. See [EHM09].

Englert:2014:NLB

[Eng14] François Englert. Nobel lecture: The BEH mechanism and its scalar boson. Reviews of Modern Physics, 86(3):843–??, July/ September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.843.

Eckmann:2019:CPP

[ERT19] Jean-Pierre Eckmann, Jacques Rougemont, and Tsvi Tlusty. Colloquium: Proteins: The physics of amorphous evolving matter. Reviews of Modern Physics, 91(3):031001–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.031001.

Efstathiou:2010:NGA

[ES10] K. Efstathiou and D. A. Sadovskií. Normalization and global analysis of perturbations of the hydrogen atom. Reviews of Modern Physics, 82(3):2099–2154, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2099;   http://rmp.aps.org/abstract/RMP/v82/i3/p2099_

Efrati:2014:RSR

[EWKK14] Efi Efrati, Zhe Wang, Amy Kolan, and Leo P. Kadanoff. Real-space renormalization in statistical mechanics. Reviews of Modern Physics, 86(2):647–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.647.

Forn-Diaz:2019:UCR

[FDLR^{+}19] P. Forn-Díaz, L. Lamata, E. Rico, J. Kono, and E. Solano. Ultrastrong coupling regimes of light-matter interaction. Reviews of Modern Physics, 91(2):025005–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.025005.

Freysoldt:2014:FPC

[FGH^{+}14] Christoph Freysoldt, Blazej Grabowski, Tilmann Hickel, Jörg Neugebauer, Georg Kresse, Anderson Janotti, and Chris G. Van de Walle. First-principles calculations for point defects in solids. Reviews of Modern Physics, 86(1):253–??, January/ March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.86.253.

Feng:2010:SSH

[FGN10] Jonathan L. Feng, Jean-François Grivaz, and Jane Nachtman. Searches for supersymmetry at high-energy colliders. Reviews of Modern Physics, 82(1):699–727, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.699;   http://rmp.aps.org/
abstract/RMP/v82/i1/p699_1.

Fodor:2012:LHM

[FH12] Zoltan Fodor and Christian Hoelbling. Light hadron masses from lattice QCD. Reviews of Modern Physics, 84(2):449–495, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.84.449;   http://rmp.aps.org/
abstract/RMP/v84/i2/p449_1.

Freer:2018:MCL

[FHKE^{+}18] Martin Freer, Hisashi Horiuchi, Yoshiko Kanada-En’yo, Dean Lee, and Ulf-G. Meißner. Microscopic clustering in light nuclei. Reviews of Modern Physics, 90(3):035004–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.035004.

Fang:2019:NPB

[FKLW19] Xiaona Fang, Karsten Kruse, Ting Lu, and Jin Wang. Nonequilibrium physics in biology. Reviews of Modern Physics, 91 (4):045004–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
045004.

Fradkin:2015:CTI

[FKT15] Eduardo Fradkin, Steven A. Kivelson, and John M. Tranquada. Colloquium: Theory of intertwined orders in high temperature superconductors. Reviews of Modern Physics, 87(2):457–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.457.

Freese:2013:AMD

[FLS13] Katherine Freese, Mariangela Lisanti, and Christopher Savage. Colloquium: Annual modulation of dark matter. Reviews of Modern Physics, 85(4):1561–1581, October 2013. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
85.1561;   http://rmp.aps.org/abstract/RMP/v85/i4/p1561_
1.

Fennel:2010:LDN

[FMBT^{+}10] Th. Fennel, K.-H. Meiwes-Broer, J. Tiggesbäumker, P.-G. Reinhard, P. M. Dinh, and E. Suraud. Laser-driven nonlinear cluster dynamics. Reviews of Modern Physics, 82(2): 1793–1842, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1793;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1793_1.

French:2010:LRI

[FPP^{+}10] Roger H. French, V. Adrian Parsegian, Rudolf Podgornik, Rick F. Rajter, Anand Jagota, Jian Luo, Dilip Asthagiri, Manoj K. Chaudhury, Yet ming Chiang, Steve Granick, Sergei Kalinin, Mehran Kardar, Roland Kjellander, David C. Langreth, Jennifer Lewis, Steve Lustig, David Wesolowski, John S. Wettlaufer, Wai-Yim Ching, Mike Finnis, Frank Houlihan, O. Anatole von Lilienfeld, Carel Jan van Oss, and Thomas Zemb. Long range interactions in nanoscale science. Reviews of Modern Physics, 82(2):1887–1944, April 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.1887;   http://rmp.aps.org/abstract/RMP/v82/i2/p1887_
1.

Fuchs:2013:QBC

[FS13] Christopher A. Fuchs and Rüdiger Schack. Quantum-Bayesian coherence. Reviews of Modern Physics, 85(4):1693–??, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.85.1693.

Frowis:2018:MQS

[FSD^{+}18] Florian Fröwis, Pavel Sekatski, Wolfgang Dür, Nicolas Gisin, and Nicolas Sangouard. Macroscopic quantum states: Measures, fragility, and implementations. Reviews of Modern Physics, 90(2):025004–??, February 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.025004.

Firstenberg:2013:CDP

[FSRD13] O. Firstenberg, M. Shuker, A. Ron, and N. Davidson. Colloquium: Coherent diffusion of polaritons in atomic media. Reviews of Modern Physics, 85(3):941–960, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.941;   http://rmp.aps.org/
abstract/RMP/v85/i3/p941_1.

Furrer:2013:MCE

[FW13] Albert Furrer and Oliver Waldmann. Magnetic cluster excitations. Reviews of Modern Physics, 85(1):367–420, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.367;   http://rmp.aps.org/
abstract/RMP/v85/i1/p367_1.

Friend:2011:MAM

[FY11] James Friend and Leslie Y. Yeo. Microscale acoustofluidics: Microfluidics driven via acoustics and ultrasonics. Reviews of Modern Physics, 83(2):647–704, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.647;
http://rmp.aps.org/abstract/RMP/v83/i2/p647_1.

Formaggio:2012:EEN

[FZ12] J. A. Formaggio and G. P. Zeller. From eV to EeV: Neutrino cross sections across energy scales. Reviews of Modern Physics, 84(3):1307–1341, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1307;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1307_1.

Georgescu:2014:QS

[GAN14] I. M. Georgescu, S. Ashhab, and Franco Nori. Quantum simulation. Reviews of Modern Physics, 86(1):153–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.153.

Guan:2013:FGO

[GBL13] Xi-Wen Guan, Murray T. Batchelor, and Chaohong Lee. Fermi gases in one dimension: From Bethe ansatz to experiments. Reviews of Modern Physics, 85(4):1633–??, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.85.1633.

Genzel:2010:GCM

[GEG10] Reinhard Genzel, Frank Eisenhauer, and Stefan Gillessen. The Galactic Center massive black hole and nuclear star cluster. Reviews of Modern Physics, 82(4):3121–3195, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.82.3121;   http://rmp.aps.org/abstract/RMP/
v82/i4/p3121_1.

Geim:2011:NLR

[Gei11] Andre K. Geim. Nobel Lecture: Random walk to graphene. Reviews of Modern Physics, 83(3):851–862, July 2011. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.83.851;   http://rmp.aps.org/abstract/RMP/
v83/i3/p851_1.

Gardner:2010:MPO

[GGG10] Jason S. Gardner, Michel J. P. Gingras, and John E. Greedan. Magnetic pyrochlore oxides. Reviews of Modern Physics, 82 (1):53–107, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.82.53;   http:/
/rmp.aps.org/abstract/RMP/v82/i1/p53_1.

Greene:2017:UFB

[GGPR17] Chris H. Greene, P. Giannakeas, and J. Pérez-Ríos. Universal few-body physics and cluster formation. Reviews of Modern Physics, 89(3):035006–??, March 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.89.035006.

Gal:2016:SNP

[GHM16] A. Gal, E. V. Hungerford, and D. J. Millener. Strangeness in nuclear physics. Reviews of Modern Physics, 88(3):035004– ??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.88.
035004.

Guo:2018:HM

[GHM^{+}18] Feng-Kun Guo, Christoph Hanhart, Ulf-G. Meißner, Qian Wang, Qiang Zhao, and Bing-Song Zou. Hadronic molecules. Reviews of Modern Physics, 90(1):015004–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.90.015004. See erratum

[GHM^{+}22].

Guo:2022:EHM

[GHM^{+}22] Feng-Kun Guo, Christoph Hanhart, Ulf-G. Meißner, Qian Wang, Qiang Zhao, and Bing-Song Zou. Erratum: Hadronic molecules [Rev. Mod. Phys. 90, 015004 (2018)]. Reviews of Modern Physics, 94(2):029901–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.029901. See [GHM^{+}18].

Gover:2019:SSS

[GIF^{+}19] A. Gover, R. Ianconescu, A. Friedman, C. Emma, N. Sudar, P. Musumeci, and C. Pellegrini. Superradiant and stimulatedsuperradiant emission of bunched electron beams. Reviews of Modern Physics, 91(3):035003–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.035003.

Giustino:2017:EPI

[Giu17] Feliciano Giustino. Electron–phonon interactions from first principles. Reviews of Modern Physics, 89(1):015003–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.015003. See erratum

[Giu19].

Giustino:2019:EEP

[Giu19] Feliciano Giustino. Erratum: Electron–phonon interactions from first principles [Rev. Mod. Phys. 89, 15003 (2017)]. Reviews of Modern Physics, 91(1):019901–??, January 2019. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.019901. See [Giu17].

Gorkov:2018:CHP

[GK18] Lev P. Gor’kov and Vladimir Z. Kresin. Colloquium: High pressure and road to room temperature superconductivity. Reviews of Modern Physics, 90(1):011001–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.011001.

Grimvall:2012:LIM

[GMKOP12] Göran Grimvall, Blanka Magyari-Köpe, Vidvuds Ozolins, and Kristin A. Persson. Lattice instabilities in metallic elements. Reviews of Modern Physics, 84(2):945–986, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.84.945;   http://rmp.aps.org/
abstract/RMP/v84/i2/p945_1.

Gull:2011:CTM

[GML^{+}11] Emanuel Gull, Andrew J. Millis, Alexander I. Lichtenstein, Alexey N. Rubtsov, Matthias Troyer, and Philipp Werner. Continuous-time Monte Carlo methods for quantum impurity models. Reviews of Modern Physics, 83(2):349–404, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.349;   http://rmp.aps.org/
abstract/RMP/v83/i2/p349_1.

Giuliani:2019:CSE

[GMN^{+}19] S. A. Giuliani, Z. Matheson, W. Nazarewicz, E. Olsen, P.- G. Reinhard, J. Sadhukhan, B. Schuetrumpf, N. Schunck, and P. Schwerdtfeger. Colloquium: Superheavy elements: Oganesson and beyond. Reviews of Modern Physics, 91(1):011001–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.91.011001.

Green:2017:MIL

[GMP17] D. R. Green, P. Meade, and M.-A. Pleier. Multiboson interactions at the LHC. Reviews of Modern Physics, 89(3):035008–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.035008.

Goldhaber:2010:PGM

[GN10] Alfred Scharff Goldhaber and Michael Martin Nieto. Photon and graviton mass limits. Reviews of Modern Physics, 82(1): 939–979, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.82.939;   http://rmp.
aps.org/abstract/RMP/v82/i1/p939_1.

Gentile:2017:OPH

[GNSW17] T. R. Gentile, P. J. Nacher, B. Saam, and T. G. Walker. Optically polarized He^{3}. Reviews of Modern Physics, 89(4):045004– ??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

www.w3.org/1998/Math/MathML.

Goerbig:2011:EPG

[Goe11] M. O. Goerbig. Electronic properties of graphene in a strong magnetic field. Reviews of Modern Physics, 83(4): 1193–1243, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1193;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1193_1.

Goldman:2014:CBP

[Gol14] Daniel I. Goldman. Colloquium: Biophysical principles of undulatory self-propulsion in granular media. Reviews of Modern Physics, 86(3):943–??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.943.

Guery-Odelin:2019:SAC

[GORK^{+}19] D. Guéry-Odelin, A. Ruschhaupt, A. Kiely, E. Torrontegui, S. Martínez-Garaot, and J. G. Muga. Shortcuts to adiabaticity: Concepts, methods, and applications. Reviews of Modern Physics, 91(4):045001–??, April 2019. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.91.045001.

Grason:2015:CGO

[Gra15] Gregory M. Grason. Colloquium: Geometry and optimal packing of twisted columns and filaments. Reviews of Modern Physics, 87(2):401–??, April/June 2015. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.87.401.

Giunti:2015:NEI

[GS15] Carlo Giunti and Alexander Studenikin. Neutrino electromagnetic interactions: A window to new physics. Reviews of Modern Physics, 87(2):531–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.531.

Giammanco:2018:STQ

[GS18] Andrea Giammanco and Reinhard Schwienhorst. Single topquark production at the Tevatron and the LHC. Reviews of Modern Physics, 90(3):035001–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.035001.

Giraud:2010:HSD

[GT10] Olivier Giraud and Koen Thas. Hearing shapes of drums: Mathematical and physical aspects of isospectrality. Reviews of Modern Physics, 82(3):2213–2255, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2213;   http://rmp.aps.org/abstract/RMP/v82/i3/p2213_
1.

Garcia-Vidal:2010:LPT

[GVMMEK10] F. J. Garcia-Vidal, L. Martin-Moreno, T. W. Ebbesen, and L. Kuipers. Light passing through subwavelength apertures. Reviews of Modern Physics, 82(1):729–787, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.729;   http://rmp.aps.org/
abstract/RMP/v82/i1/p729_1.

Gu:2018:CPT

[GWY^{+}18] Xiaokun Gu, Yujie Wei, Xiaobo Yin, Baowen Li, and Ronggui Yang. Colloquium: Phononic thermal properties of twodimensional materials. Reviews of Modern Physics, 90(4): 041002–??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.90.
041002.

Gribakin:2010:PMI

[GYS10] G. F. Gribakin, J. A. Young, and C. M. Surko. Positronmolecule interactions: Resonant attachment, annihilation, and bound states. Reviews of Modern Physics, 82(3):2557– 2607, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2557;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2557_1.

Haldane:2017:NLT

[Hal17] F. Duncan M. Haldane. Nobel lecture: Topological quantum matter. Reviews of Modern Physics, 89(4):040502–??, October/ December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.89.040502.

Haroche:2013:NLC

[Har13] Serge Haroche. Nobel Lecture: Controlling photons in a box and exploring the quantum to classical boundary. Reviews of Modern Physics, 85(3):1083–1102, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.85.
1083;   http://rmp.aps.org/abstract/RMP/v85/i3/p1083_
1.

Harlow:2016:JLB

[Har16] D. Harlow. Jerusalem lectures on black holes and quantum information. Reviews of Modern Physics, 88(1):015002–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.015002.

Herrero-Collantes:2017:QRN

[HCGE17] Miguel Herrero-Collantes and Juan Carlos Garcia-Escartin. Quantum random number generators. Reviews of Modern Physics, 89(1):015004:1–015004:48, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.015004.

Hell:2015:NLN

[Hel15] Stefan W. Hell. Nobel lecture: Nanoscopy with freely propagating light. Reviews of Modern Physics, 87(4):1169–??, October/ December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.1169.

Hornberger:2012:CQI

[HGH^{+}12] Klaus Hornberger, Stefan Gerlich, Philipp Haslinger, Stefan Nimmrichter, and Markus Arndt. Colloquium: Quantum interference of clusters and molecules. Reviews of Modern Physics, 84(1):157–173, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.84.157;   http:
//rmp.aps.org/abstract/RMP/v84/i1/p157_1.

Hayano:2010:HPN

[HH10] Ryugo S. Hayano and Tetsuo Hatsuda. Hadron properties in the nuclear medium. Reviews of Modern Physics, 82(4): 2949–2990, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2949;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p2949_1.

Hansson:2017:QHP

[HHSV17] T. H. Hansson, M. Hermanns, S. H. Simon, and S. F. Viefers. Quantum Hall physics: Hierarchies and conformal field theory techniques. Reviews of Modern Physics, 89(2):025005–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.025005.

Higgs:2014:NLE

[Hig14] Peter W. Higgs. Nobel lecture: Evading the Goldstone theorem. Reviews of Modern Physics, 86(3):851–??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.851.

Hinterbichler:2012:TAM

[Hin12] Kurt Hinterbichler. Theoretical aspects of massive gravity. Reviews of Modern Physics, 84(2):671–710, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.84.671;   http://rmp.aps.org/
abstract/RMP/v84/i2/p671_1.

Hasan:2010:CTI

[HK10] M. Z. Hasan and C. L. Kane. Colloquium: Topological insulators. Reviews of Modern Physics, 82(4):3045– 3067, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.3045;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p3045_1.

Hartnett:2011:CCA

[HL11] John G. Hartnett and Andre N. Luiten. Colloquium: Comparison of astrophysical and terrestrial frequency standards. Reviews of Modern Physics, 83(1):1–9, January 2011. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
83.1;   http://rmp.aps.org/abstract/RMP/v83/i1/p1_1.

Hen:2017:NNC

[HMPW17] Or Hen, Gerald A. Miller, Eli Piasetzky, and Lawrence B. Weinstein. Nucleon–nucleon correlations, short-lived excitations, and the quarks within. Reviews of Modern Physics, 89(4):045002–??, October/December 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.89.045002.

Hammer:2013:CTB

[HNS13] Hans-Werner Hammer, Andreas Nogga, and Achim Schwenk. Colloquium: Three-body forces: From cold atoms to nuclei. Reviews of Modern Physics, 85(1):197–217, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.197;   http://rmp.aps.org/
abstract/RMP/v85/i1/p197_1.

Hobbs:2012:TSE

[HNW12] John D. Hobbs, Mark S. Neubauer, and Scott Willenbrock. Tests of the standard electroweak model at the energy frontier. Reviews of Modern Physics, 84(4):1477– 1526, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1477;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1477_1.

Hohenberg:2010:CIC

[Hoh10] P. C. Hohenberg. Colloquium: An introduction to consistent quantum theory. Reviews of Modern Physics, 82(4):

2835–2844, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2835;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p2835_1.

Holt:2010:NPD

[HR10] Roy J. Holt and Craig D. Roberts. Nucleon and pion distribution functions in the valence region. Reviews of Modern Physics, 82(4):2991–3044, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2991;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p2991_1.

Hayen:2018:HPA

[HSB^{+}18] Leendert Hayen, Nathal Severijns, Kazimierz Bodek, Dagmara Rozpedzik, and Xavier Mougeot. High precision analytical description of the allowed β spectrum shape. Reviews of Modern Physics, 90(1):015008–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://www.w3.org/1998/Math/MathML.

Hammerer:2010:QIB

[HSP10] Klemens Hammerer, Anders S. Sørensen, and Eugene S. Polzik. Quantum interface between light and atomic ensembles. Reviews of Modern Physics, 82(2):1041–1093, April 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.1041;   http://rmp.aps.org/abstract/RMP/v82/i2/p1041_
1.

Hemsing:2014:BDL

[HSXZ14] Erik Hemsing, Gennady Stupakov, Dao Xiang, and Alexander Zholents. Beam by design: Laser manipulation of electrons in modern accelerators. Reviews of Modern Physics, 86(3):897– ??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.86.
897.

etal:2017:IIP

[HTB^{+}17] Frances Hellman Axel Hoffmann, Yaroslav Tserkovnyak, Geoffrey S. D. Beach, Eric E. Fullerton, Chris Leighton, Allan H. MacDonald, Daniel C. Ralph, Dario A. Arena, Hermann A. Dürr, Peter Fischer, Julie Grollier, Joseph P. Heremans, Tomas Jungwirth, Alexey V. Kimel, Bert Koopmans, Ilya N. Krivorotov, Steven J. May, Amanda K. Petford-Long, James M. Rondinelli, Nitin Samarth, Ivan K. Schuller, Andrei N. Slavin, Mark D. Stiles, Oleg Tchernyshyov, André Thiaville, and Barry L. Zink. Interface-induced phenomena in magnetism. Reviews of Modern Physics, 89(2):025006–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.025006.

Heyde:2010:MDE

[HvNCR10] Kris Heyde, Peter von Neumann-Cosel, and Achim Richter. Magnetic dipole excitations in nuclei: Elementary modes of nucleonic motion. Reviews of Modern Physics, 82(3):2365– 2419, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2365;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2365_1.

Heyde:2011:PNS

[HW11a] Kris Heyde and John L. Wood. Publisher’s note: Shape coexistence in atomic nuclei [Rev. Mod. Phys. 83, 1467 (2011)]. Reviews of Modern Physics, 83(4):1655, October 2011. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.83.1655;   http://rmp.aps.org/abstract/RMP/
v83/i4/p1655_1. See [HW11b].

Heyde:2011:SCA

[HW11b] Kris Heyde and John L. Wood. Shape coexistence in atomic nuclei. Reviews of Modern Physics, 83(4):1467– 1521, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1467;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1467_1. See note

[HW11a].

Imambekov:2012:ODQ

[ISG12] Adilet Imambekov, Thomas L. Schmidt, and Leonid I. Glazman. One-dimensional quantum liquids: Beyond the Luttinger liquid paradigm. Reviews of Modern Physics, 84(3): 1253–1306, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1253;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1253_1.

Jensen:2016:SFS

[JBSB^{+}16] K. H. Jensen, K. Berg-Sørensen, H. Bruus, N. M. Holbrook, J. Liesche, A. Schulz, M. A. Zwieniecki, and T. Bohr. Sap flow and sugar transport in plants. Reviews of Modern Physics, 88(3):035007–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.035007.

Jakli:2018:PLC

[JLS18] Antal Jákli, Oleg D. Lavrentovich, and Jonathan V. Selinger. Physics of liquid crystals of bent-shaped molecules. Reviews of Modern Physics, 90(4):045004–??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.045004.

Jang:2018:DEN

[JM18] Seogjoo J. Jang and Benedetta Mennucci. Delocalized excitons in natural light-harvesting complexes. Reviews of Modern Physics, 90(3):035003–??, March 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.035003.

Jones:2015:DFT

[Jon15] R. O. Jones. Density functional theory: Its origins, rise to prominence, and future. Reviews of Modern Physics, 87(3):897– 923, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.87.
897.

Jain:2017:NPQ

[JS17] Sudhir Ranjan Jain and Rhine Samajdar. Nodal portraits of quantum billiards: Domains, lines, and statistics. Reviews of Modern Physics, 89(4):045005–??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.045005.

Jungwirth:2014:SDP

[JWN^{+}14] T. Jungwirth, J. Wunderlich, V. Novák, K. Olejník, B. L. Gallagher, R. P. Campion, K. W. Edmonds, A. W. Rushforth, A. J. Ferguson, and P. Nemec. Spin-dependent phenomena and device concepts explored in (ga,mn)as. Reviews of Modern Physics, 86(3):855–??, July/September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.855.

Kikuchi:2012:SST

[KA12] M. Kikuchi and M. Azumi. Steady-state tokamak research: Core physics. Reviews of Modern Physics, 84(4):1807– 1854, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1807;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1807_1.

Kashlinsky:2018:LCN

[KAAB^{+}18] A. Kashlinsky, R. G. Arendt, F. Atrio-Barandela, N. Cappelluti, A. Ferrara, and G. Hasinger. Looking at cosmic near-infrared background radiation anisotropies. Reviews of Modern Physics, 90(2):025006–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.025006.

Kamien:2019:ERL

[KAB^{+}19] Randall D. Kamien, Hiroaki Aihara, Dietrich Belitz, Debbie Brodbar, A. H Castro Neto, Margaret S. Cheung, William D. Collins, Marjolein Dijkstra, David DiVincenzo, Paul D. Grannis, Arthur F. Hebard, Vicky Kalogera, Igor Klebanov, Wim Leemans, Klaus Mølmer, Witold Nazarewicz, Pierre Ramond, Roxanne Springer, Anthony F. Starace, and Friedel Thielemann. Editorial: RMP: Looking forward. Reviews of Modern Physics, 91(3):030001–??, March 2019. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.91.030001.

Korolev:2010:GDE

[KAHN10] K. S. Korolev, Mikkel Avlund, Oskar Hallatschek, and David R. Nelson. Genetic demixing and evolution in linear stepping stone models. Reviews of Modern Physics, 82(2):1691– 1718, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1691;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1691_1.

Kajita:2016:NLD

[Kaj16] Takaaki Kajita. Nobel Lecture: Discovery of atmospheric neutrino oscillations. Reviews of Modern Physics, 88(3):030501– ??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.88.
030501.

Kamien:2019:ESC

[Kam19] Randall Kamien. Editorial: Synthesizing current research succinctly and elegantly. Reviews of Modern Physics, 91(1):010001– ??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.91.010001.

Kao:2010:NLS

[Kao10] Charles K. Kao. Nobel Lecture: Sand from centuries past: Send future voices fast. Reviews of Modern Physics, 82(3): 2299–2303, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2299;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2299_1.

Kim:2010:ASC

[KC10] Jihn E. Kim and Gianpaolo Carosi. Axions and the strong CP problem. Reviews of Modern Physics, 82(1):557–601, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.557;   http://rmp.aps.org/
abstract/RMP/v82/i1/p557_1. See erratum [KC19].

Kim:2019:EAS

[KC19] Jihn E. Kim and Gianpaolo Carosi. Erratum: Axions and the strong CP problem [Rev. Mod. Phys. 82, 557 (2010)]. Reviews of Modern Physics, 91(4):049902–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://www.w3.org/1998/Math/MathML. See

[KC10].

Kappeler:2011:PNP

[KGBA11] F. Käppeler, R. Gallino, S. Bisterzo, and Wako Aoki. The s process: Nuclear physics, stellar models, and observations. Reviews of Modern Physics, 83(1):157–193, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.157;   http://rmp.aps.org/
abstract/RMP/v83/i1/p157_1.

Kawamura:2012:SPF

[KHK^{+}12] Hikaru Kawamura, Takahiro Hatano, Naoyuki Kato, Soumyajyoti Biswas, and Bikas K. Chakrabarti. Statistical physics of fracture, friction, and earthquakes. Reviews of Modern Physics, 84(2):839–884, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.84.839;   http://rmp.
aps.org/abstract/RMP/v84/i2/p839_1.

Kirilyuk:2010:UOM

[KKR10] Andrei Kirilyuk, Alexey V. Kimel, and Theo Rasing. Ultrafast optical manipulation of magnetic order. Reviews of Modern Physics, 82(3):2731–2784, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2731;   http://rmp.aps.org/abstract/RMP/v82/i3/p2731_

1. See erratum [KKR16].

Kirilyuk:2016:EUO

[KKR16] Andrei Kirilyuk, Alexey V. Kimel, and Theo Rasing. Erratum: Ultrafast optical manipulation of magnetic order [Rev. Mod. Phys. 82, 2731 (2010)]. Reviews of Modern Physics, 88(3):039904–??, July/September 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.88.039904. See [KKR10].

Koster:2012:SPP

[KKS^{+}12] Gertjan Koster, Lior Klein, Wolter Siemons, Guus Rijnders, J. Steven Dodge, Chang-Beom Eom, Dave H. A. Blank, and Malcolm R. Beasley. Structure, physical properties, and applications of SrRuO_{3} thin films. Reviews of Modern Physics, 84(1): 253–298, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.84.253;   http://rmp.
aps.org/abstract/RMP/v84/i1/p253_1.

Kapitulnik:2019:CAM

[KKS19] Aharon Kapitulnik, Steven A. Kivelson, and Boris Spivak. Colloquium: Anomalous metals: Failed superconductors. Reviews of Modern Physics, 91(1):011002–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.011002.

Kruchinin:2018:CSF

[KKY18] Stanislav Yu. Kruchinin, Ferenc Krausz, and Vladislav S. Yakovlev. Colloquium: Strong-field phenomena in periodic systems. Reviews of Modern Physics, 90(2):021002–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.021002.

Kartashov:2011:PNS

[KMT11a] Yaroslav V. Kartashov, Boris A. Malomed, and Lluis Torner. Publisher’s note: Solitons in nonlinear lattices [rev. mod. phys. 83, 247 (2011)]. Reviews of Modern Physics, 83(2):405, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.405;   http://rmp.aps.org/
abstract/RMP/v83/i2/p405_1. See [KMT11b].

Kartashov:2011:SNL

[KMT11b] Yaroslav V. Kartashov, Boris A. Malomed, and Lluis Torner. Solitons in nonlinear lattices. Reviews of Modern Physics, 83(1): 247–305, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.83.247;   http://rmp.
aps.org/abstract/RMP/v83/i1/p247_1. See note [KMT11a].

Kogler:2019:JSL

[KNS^{+}19] Roman Kogler, Benjamin Nachman, Alexander Schmidt, Lily Asquith, Emma Winkels, Mario Campanelli, Chris Delitzsch, Philip Harris, Andreas Hinzmann, Deepak Kar, Christine McLean, Justin Pilot, Yuta Takahashi, Nhan Tran, Caterina Vernieri, and Marcel Vos. Jet substructure at the Large Hadron Collider. Reviews of Modern Physics, 91(4):045003–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.045003.

Kosterlitz:2017:NLT

[Kos17] John Michael Kosterlitz. Nobel lecture: Topological defects and phase transitions. Reviews of Modern Physics, 89(4):040501–??, October/December 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.89.040501.

Keim:2019:MFM

[KPZ^{+}19] Nathan C. Keim, Joseph D. Paulsen, Zorana Zeravcic, Srikanth Sastry, and Sidney R. Nagel. Memory formation in matter. Reviews of Modern Physics, 91(3):035002–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.035002.

Klempt:2010:BS

[KR10] Eberhard Klempt and Jean-Marc Richard. Baryon spectroscopy. Reviews of Modern Physics, 82(2):1095–1153, April 2010. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.82.1095;   http://rmp.aps.org/abstract/RMP/
v82/i2/p1095_1.

Kostelecky:2011:DTL

[KR11] V. Alan Kostelecký and Neil Russell. Data tables for Lorentz and CPT violation. Reviews of Modern Physics, 83(1):11– 31, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.11;   http://
rmp.aps.org/abstract/RMP/v83/i1/p11_1.

Kozlov:2018:HCI

[KSLUS18] M. G. Kozlov, M. S. Safronova, J. R. Crespo López-Urrutia, and P. O. Schmidt. Highly charged ions: Optical clocks and applications in fundamental physics. Reviews of Modern Physics, 90(4):045005–??, April 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.045005.

Kirkpatrick:2015:CRF

[KT15] T. R. Kirkpatrick and D. Thirumalai. Colloquium: Random first order transition theory concepts in biology and physics. Reviews of Modern Physics, 87(1):183–??, January/ March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.87.183.

Kotov:2012:EEI

[KUP^{+}12] Valeri N. Kotov, Bruno Uchoa, Vitor M. Pereira, F. Guinea, and A. H. Castro Neto. Electron-electron interactions in graphene: Current status and perspectives. Reviews of Modern Physics, 84(3):1067–1125, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1067;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1067_1.

Kosztin:2012:CMD

[KVNF12] Ioan Kosztin, Gordana Vunjak-Novakovic, and Gabor Forgacs. Colloquium: Modeling the dynamics of multicellular systems: Application to tissue engineering. Reviews of Modern Physics, 84(4):1791–1805, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1791;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1791_1.

Konotop:2016:NWP

[KYZ16] Vladimir V. Konotop, Jianke Yang, and Dmitry A. Zezyulin. Nonlinear waves in PT -symmetric systems. Reviews of Modern Physics, 88(3):035002–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.035002.

Konoplya:2011:QMB

[KZ11] R. A. Konoplya and Alexander Zhidenko. Quasinormal modes of black holes: From astrophysics to string theory. Reviews of Modern Physics, 83(3):793–836, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.793;
http://rmp.aps.org/abstract/RMP/v83/i3/p793_1.

Lalli:2014:NSJ

[Lal14] Roberto Lalli. A new scientific journal takes the scene: The birth of Reviews of Modern Physics. Annalen der Physik 8 (Berlin, Germany), 526(9–10):A83–A87, October 2014. CODEN AN-

PYA2. ISSN 0003-3804. URL http://onlinelibrary.wiley.
com/enhanced/doi/10.1002/andp.201400810/.

Liu:2016:CPC

[LB16] Yang-Yu Liu and Albert-László Barabási. Control principles of complex systems. Reviews of Modern Physics, 88(3):035006– ??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.88.
035006.

Linder:2019:OFS

[LB19] Jacob Linder and Alexander V. Balatsky. Odd-frequency superconductivity. Reviews of Modern Physics, 91(4):045005–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.91.045005.

Lloyd:2017:EVB

[LBTY17] S. M. Lloyd, M. Babiker, G. Thirunavukkarasu, and J. Yuan. Electron vortices: Beams with orbital angular momentum. Reviews of Modern Physics, 89(3):035004–??, March 2017. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.035004.

Ludlow:2015:OAC

[LBY^{+}15] Andrew D. Ludlow, Martin M. Boyd, Jun Ye, E. Peik, and P. O. Schmidt. Optical atomic clocks. Reviews of Modern Physics, 87(2):637–??, April/June 2015. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.87.637.

Lebedev:2019:EAR

[LFR19] S. V. Lebedev, A. Frank, and D. D. Ryutov. Exploring astrophysics-relevant magnetohydrodynamics with pulsedpower laboratory facilities. Reviews of Modern Physics, 91(2): 025002–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
025002.

Laird:2015:QTC

[LKS^{+}15] Edward A. Laird, Ferdinand Kuemmeth, Gary A. Steele, Kasper Grove-Rasmussen, Jesper Nygård, Karsten Flensberg, and Leo P. Kouwenhoven. Quantum transport in carbon nanotubes. Reviews of Modern Physics, 87(3):703–??, July/ September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.703.

Lingam:2019:CPC

[LL19] Manasvi Lingam and Abraham Loeb. Colloquium: Physical constraints for the evolution of life on exoplanets. Reviews of Modern Physics, 91(2):021002–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.021002.

LaNave:2019:CFE

[LLP19] Gabriele La Nave, Kridsanaphong Limtragool, and Philip W. Phillips. Colloquium: Fractional electromagnetism in quantum matter and high-energy physics. Reviews of Modern Physics, 91(2):021003–??, February 2019. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.91.021003.

Lee:2018:DRC

[LLW18] Sehwook Lee, Michele Livan, and Richard Wigmans. Dualreadout calorimetry. Reviews of Modern Physics, 90(2):025002– ??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.025002.

Lu:2013:LPN

[LMD^{+}13] Z.-T. Lu, P. Mueller, G. W. F. Drake, W. Nörtershäuser, Steven C. Pieper, and Z.-C. Yan. Colloquium: Laser probing of neutron-rich nuclei in light atoms. Reviews of Modern Physics, 85(4):1383–1400, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1383;   http:
//rmp.aps.org/abstract/RMP/v85/i4/p1383_1.

Lodahl:2015:ISP

[LMS15] Peter Lodahl, Sahand Mahmoodian, and Søren Stobbe. Interfacing single photons and single quantum dots with photonic nanostructures. Reviews of Modern Physics, 87(2):347–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.87.347.

Londergan:2010:CSP

[LPT10] J. T. Londergan, J. C. Peng, and A. W. Thomas. Charge symmetry at the partonic level. Reviews of Modern Physics, 82 (3):2009–2052, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2009;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2009_1.

Li:2012:CPM

[LRW^{+}12] Nianbei Li, Jie Ren, Lei Wang, Gang Zhang, Peter Hänggi, and Baowen Li. Colloquium: Phononics: Manipulating heat flow with electronic analogs and beyond. Reviews of Modern Physics, 84(3):1045–1066, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1045;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1045_1.

Lapine:2014:CNM

[LSK14] Mikhail Lapine, Ilya V. Shadrivov, and Yuri S. Kivshar. Colloquium: Nonlinear metamaterials. Reviews of Modern Physics, 86(3):1093–??, January/March 2014. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.86.1093.

Letessier-Selvon:2011:UEC

[LSS11] Antoine Letessier-Selvon and Todor Stanev. Ultrahigh energy cosmic rays. Reviews of Modern Physics, 83(3):907–942, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.907;   http://rmp.aps.org/
abstract/RMP/v83/i3/p907_1.

Lohse:2015:SNN

[LZ15] Detlef Lohse and Xuehua Zhang. Surface nanobubbles and nanodroplets. Reviews of Modern Physics, 87(3):981–??, July/ September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.87.981.

Modi:2012:CQB

[MBC^{+}12] Kavan Modi, Aharon Brodutch, Hugo Cable, Tomasz Paterek, and Vlatko Vedral. The classical-quantum boundary for correlations: Discord and related measures. Reviews of Modern Physics, 84(4):1655–1707, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.84.
1655;   http://rmp.aps.org/abstract/RMP/v84/i4/p1655_
1.

Mino:2018:MCS

[MBSR^{+}18] Lorenzo Mino, Elisa Borfecchia, Jaime Segura-Ruiz, Cinzia Giannini, Gema Martinez-Criado, and Carlo Lamberti. Materials characterization by synchrotron X-ray microprobes and nanoprobes. Reviews of Modern Physics, 90(2):025007–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.025007.

McDonald:2016:NLS

[McD16] Arthur B. McDonald. Nobel Lecture: The Sudbury Neutrino Observatory: Observation of flavor change for solar neutrinos. Reviews of Modern Physics, 88(3):030502–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.030502.

Mao:2018:SLG

[MCD^{+}18] Ho-Kwang Mao, Xiao-Jia Chen, Yang Ding, Bing Li, and Lin Wang. Solids, liquids, and gases under high pressure. Reviews of Modern Physics, 90(1):015007–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.015007.

Maciolek:2018:CBC

[MD18] Anna Maciolek and Siegfried Dietrich. Collective behavior of colloids due to critical Casimir interactions. Reviews of Modern Physics, 90(4):045001–??, April 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.045001.

Mermin:1993:HVT

[Mer93] N. David Mermin. Hidden variables and the two theorems of John Bell. Reviews of Modern Physics, 65(3):803–815, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.65.803;   http://rmp.aps.org/
abstract/RMP/v65/i3/p803_1. See errata [Mer16, Mer17].

Mermin:2016:EHV

[Mer16] N. David Mermin. Erratum: Hidden variables and the two theorems of John Bell [Rev. Mod. Phys. 65, 803 (1993)]. Reviews of Modern Physics, 88(3):039902–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.039902. See [Mer93].

Mermin:2017:EHV

[Mer17] N. David Mermin. Erratum: Hidden variables and the two theorems of John Bell [Rev. Mod. Phys. 65, 803 (1993)]. Reviews of Modern Physics, 89(4):049901–??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.049901. See [Mer93].

Metcalf:2017:CSO

[Met17] Harold Metcalf. Colloquium: Strong optical forces on atoms in multifrequency light. Reviews of Modern Physics, 89(4):041001– ??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.89.
041001.

Meystre:2017:EAE

[Mey17] Pierre Meystre. Editorial: From the APS Editor in Chief. Reviews of Modern Physics, 89(1):010001–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.010001.

Miroshnichenko:2010:FRN

[MFK10] Andrey E. Miroshnichenko, Sergej Flach, and Yuri S. Kivshar. Fano resonances in nanoscale structures. Reviews of Modern Physics, 82(3):2257–2298, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2257;   http://rmp.aps.org/abstract/RMP/v82/i3/p2257_
1.

Muhlbauer:2019:MSA

[MHP^{+}19] Sebastian Mühlbauer, Dirk Honecker, Élio A. Périgo, Frank Bergner, Sabrina Disch, André Heinemann, Sergey Erokhin, Dmitry Berkov, Chris Leighton, Morten Ring Eskildsen, and Andreas Michels. Magnetic small-angle neutron scattering. Reviews of Modern Physics, 91(1):015004–??, January 2019. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.015004.

Marchetti:2013:HSA

[MJR^{+}13] M. C. Marchetti, J. F. Joanny, S. Ramaswamy, T. B. Liverpool, J. Prost, Madan Rao, and R. Aditi Simha. Hydrodynamics of soft active matter. Reviews of Modern Physics, 85 (3):1143–1189, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1143;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1143_1.

Maher:2015:CYM

[MJT15] Simon Maher, Fred P. M. Jjunju, and Stephen Taylor. Colloquium: 100 years of mass spectrometry: Perspectives and future trends. Reviews of Modern Physics, 87(1):113–??, January/ March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.87.113.

Maeder:2012:RMS

[MM12] André Maeder and Georges Meynet. Rotating massive stars: From first stars to gamma ray bursts. Reviews of Modern Physics, 84(1):25–63, January 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.25;   http://rmp.aps.org/abstract/RMP/v84/i1/p25_
1.

McMahon:2012:PHH

[MMPC12] Jeffrey M. McMahon, Miguel A. Morales, Carlo Pierleoni, and David M. Ceperley. The properties of hydrogen and helium under extreme conditions. Reviews of Modern Physics, 84(4): 1607–1653, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1607;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1607_1.

Marzari:2012:MLW

[MMY^{+}12] Nicola Marzari, Arash A. Mostofi, Jonathan R. Yates, Ivo Souza, and David Vanderbilt. Maximally localized Wannier functions: Theory and applications. Reviews of Modern Physics, 84(4):1419–1475, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1419;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1419_1.

Mydosh:2011:CHO

[MO11] J. A. Mydosh and P. M. Oppeneer. Colloquium: Hidden order, superconductivity, and magnetism: The unsolved case of URu_{2}Si_{2}. Reviews of Modern Physics, 83(4):1301– 1322, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1301;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1301_1.

Moerner:2015:NLS

[Moe15] W. E. (William E.) Moerner. Nobel lecture: Singlemolecule spectroscopy, imaging, and photocontrol: Foundations for super-resolution microscopy. Reviews of Modern Physics, 87(4):1183–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1183.

Mourou:2019:NLE

[Mou19] Gerard Mourou. Nobel Lecture: Extreme light physics and application. Reviews of Modern Physics, 91(3):030501–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.030501.

Marconnet:2013:TCP

[MPG13] Amy M. Marconnet, Matthew A. Panzer, and Kenneth E. Goodson. Thermal conduction phenomena in carbon nanotubes and related nanostructured materials. Reviews of Modern Physics, 85(3):1295–1326, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1295;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1295_1.

Misbah:2010:CSE

[MPLS10] Chaouqi Misbah, Olivier Pierre-Louis, and Yukio Saito. Crystal surfaces in and out of equilibrium: a modern view. Reviews of Modern Physics, 82(1):981–1040, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.981;   http://rmp.aps.org/
abstract/RMP/v82/i1/p981_1.

Mitchell:2010:RMC

[MRW10] G. E. Mitchell, A. Richter, and H. A. Weidenmüller. Random matrices and chaos in nuclear physics: Nuclear reactions. Reviews of Modern Physics, 82(4):2845–2901, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.82.2845;   http://rmp.aps.org/abstract/RMP/
v82/i4/p2845_1.

Matagne:2015:BRL

[MS15] N. Matagne and Fl. Stancu. Baryon resonances in large N_{c} QCD. Reviews of Modern Physics, 87(1):211–??, January/ March 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.87.211.

Meunier:2016:PPL

[MSBD16] V. Meunier, A. G. Souza Filho, E. B. Barros, and M. S. Dresselhaus. Physical properties of low-dimensional sp^{2}-based carbon nanostructures. Reviews of Modern Physics, 88(2):025005–??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.w3.org/

1998/Math/MathML.

Metzner:2012:FRG

[MSH^{+}12] Walter Metzner, Manfred Salmhofer, Carsten Honerkamp, Volker Meden, and Kurt Schönhammer. Functional renormalization group approach to correlated fermion systems. Reviews of Modern Physics, 84(1):299–352, January 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.299;   http://rmp.aps.org/abstract/RMP/v84/i1/p299_
1.

Markos:2017:HPC

[MTA^{+}17] Christos Markos, John C. Travers, Amir Abdolvand, Benjamin J. Eggleton, and Ole Bang. Hybrid photonic-crystal fiber. Reviews of Modern Physics, 89(4):045003–??, October/ December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.89.045003.

Mohr:2012:CRV

[MTN12] Peter J. Mohr, Barry N. Taylor, and David B. Newell. CODATA recommended values of the fundamental physical constants: 2010. Reviews of Modern Physics, 84(4):1527– 1605, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1527;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1527_1.

Munoz:2018:CCD

[Muñ18] Miguel A. Muñoz. Colloquium: Criticality and dynamical scaling in living systems. Reviews of Modern Physics, 90(3):031001– ??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.031001.

Nagel:2017:ESM

[Nag17] Sidney R. Nagel. Experimental soft-matter science. Reviews of Modern Physics, 89(2):025002–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.025002.

Nakamura:2015:NLB

[Nak15] Shuji Nakamura. Nobel lecture: Background story of the invention of efficient blue InGaN light emitting diodes. Reviews of Modern Physics, 87(4):1139–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.87.1139.

Nambu:2009:NLS

[Nam09] Yoichiro Nambu. Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization. Reviews of Modern Physics, 81(3):1015–1018, July 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1015;   http://rmp.aps.org/abstract/RMP/v81/i3/p1015_

1. See note [Nam10].

Nambu:2010:PNN

[Nam10] Yoichiro Nambu. Publisher’s note: Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization [Rev. Mod. Phys. 81, 1015 (2009)]. Reviews of Modern Physics, 82(4):3199, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
3199;   http://rmp.aps.org/abstract/RMP/v82/i4/p3199_

1. See [Nam09].

etal:2014:DLA

[NBD^{+}14] R. Joel England Robert J. Noble, Karl Bane, David H. Dowell, Cho-Kuen Ng, James E. Spencer, Sami Tantawi, Ziran Wu, Robert L. Byer, Edgar Peralta, Ken Soong, Chia-Ming Chang, Behnam Montazeri, Stephen J. Wolf, Benjamin Cowan, Jay Dawson, Wei Gai, Peter Hommelhoff, Yen-Chieh Huang, Chunguang Jing, Christopher McGuinness, Robert B. Palmer, Brian Naranjo, James Rosenzweig, Gil Travish, Amit Mizrahi, Levi Schachter, Christopher Sears, Gregory R. Werner, and Rodney B. Yoder. Dielectric laser accelerators. Reviews of Modern Physics, 86(4):1337–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.1337.

Nicolas:2018:DFA

[NFMB18] Alexandre Nicolas, Ezequiel E. Ferrero, Kirsten Martens, and Jean-Louis Barrat. Deformation and flow of amorphous solids: Insights from elastoplastic models. Reviews of Modern Physics, 90(4):045006–??, April 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.045006.

Nishioka:2018:EEH

[Nis18] Tatsuma Nishioka. Entanglement entropy: Holography and renormalization group. Reviews of Modern Physics, 90(3): 035007–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.90.
035007.

Nation:2012:CSU

[NJBN12] P. D. Nation, J. R. Johansson, M. P. Blencowe, and Franco Nori. Colloquium: Stimulating uncertainty: Amplifying the quantum vacuum with superconducting circuits. Reviews of Modern Physics, 84(1):1–24, January 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.1;   http://rmp.aps.org/abstract/RMP/v84/i1/p1_1.

Narozhny:2016:CD

[NL16] B. N. Narozhny and A. Levchenko. Coulomb drag. Reviews of Modern Physics, 88(2):025003–??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.025003.

Nisoli:2013:ASI

[NMS13] Cristiano Nisoli, Roderich Moessner, and Peter Schiffer. Colloquium: Artificial spin ice: Designing and imaging magnetic frustration. Reviews of Modern Physics, 85(4):1473– 1490, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1473;   http:
//rmp.aps.org/abstract/RMP/v85/i4/p1473_1.

Novoselov:2011:NLG

[Nov11] K. S. Novoselov. Nobel Lecture: Graphene: Materials in the Flatland. Reviews of Modern Physics, 83(3):837–849, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.837;   http://rmp.aps.org/
abstract/RMP/v83/i3/p837_1.

Neher:2011:SGE

[NS11] Richard A. Neher and Boris I. Shraiman. Statistical genetics and evolution of quantitative traits. Reviews of Modern Physics, 83 (4):1283–1300, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1283;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1283_1.

Nagaosa:2010:AHE

[NSO^{+}10] Naoto Nagaosa, Jairo Sinova, Shigeki Onoda, A. H. MacDonald, and N. P. Ong. Anomalous Hall effect. Reviews of Modern Physics, 82(2):1539–1592, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1539;   http://rmp.aps.org/abstract/RMP/v82/i2/p1539_

Nussinov:2015:CMT

[NvdB15] Zohar Nussinov and Jeroen van den Brink. Compass models: Theory and physical motivations. Reviews of Modern Physics, 87(1):1–??, January/March 2015. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.87.1.

Newman:2014:HFS

[NW14] Paul R. Newman and Matthew Wing. The hadronic final state at HERA. Reviews of Modern Physics, 86(3):1037–??, July/ September 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.1037.

Ortiz-Ambriz:2019:CIR

[OANR^{+}19] Antonio Ortiz-Ambriz, Cristiano Nisoli, Charles Reichhardt, Cynthia J. O. Reichhardt, and Pietro Tierno. Colloquium: Ice rule and emergent frustration in particle ice and beyond. Reviews of Modern Physics, 91(4):041003–??, April 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.041003.

Oka:2014:SPQ

[OBC^{+}14] Hirofumi Oka, Oleg O. Brovko, Marco Corbetta, Valeri S. Stepanyuk, Dirk Sander, and Jürgen Kirschner. Spin-polarized quantum confinement in nanostructures: Scanning tunneling microscopy. Reviews of Modern Physics, 86(4):1127–??, April/ June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.86.1127.

Ostrikov:2016:CNG

[OBN16] Kostya (Ken) Ostrikov, Farhat Beg, and Andrew Ng. Colloquium: Nanoplasmas generated by intense radiation. Reviews of Modern Physics, 88(1):011001–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.011001.

Ozawa:2019:TP

[OPA^{+}19] Tomoki Ozawa, Hannah M. Price, Alberto Amo, Nathan Goldman, Mohammad Hafezi, Ling Lu, Mikael C. Rechtsman, David Schuster, Jonathan Simon, Oded Zilberberg, and Iacopo Carusotto. Topological photonics. Reviews of Modern Physics, 91(1):015006–??, January 2019. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.91.015006.

Olsen:2018:NHM

[OSZ18] Stephen Lars Olsen, Tomasz Skwarnicki, and Daria Zieminska. Nonstandard heavy mesons and baryons: Experimental evidence. Reviews of Modern Physics, 90(1):015003–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.015003.

Prantzos:2011:KEP

[PBB^{+}11] N. Prantzos, C. Boehm, A. M. Bykov, R. Diehl, K. Ferrière, N. Guessoum, P. Jean, J. Knoedlseder, A. Marcowith, I. V. Moskalenko, A. Strong, and G. Weidenspointner. The 511 keV emission from positron annihilation in the Galaxy. Reviews of Modern Physics, 83(3):1001–1056, July 2011. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
83.1001;   http://rmp.aps.org/abstract/RMP/v83/i3/p1001_
1.

Pan:2012:MEI

[PCL^{+}12] Jian-Wei Pan, Zeng-Bing Chen, Chao-Yang Lu, Harald Weinfurter, Anton Zeilinger, and Marek Zukowski. Multiphoton entanglement and interferometry. Reviews of Modern Physics, 84(2):777–838, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.84.777;   http://rmp.
aps.org/abstract/RMP/v84/i2/p777_1.

Peres:2010:CTP

[Per10] N. M. R. Peres. Colloquium: The transport properties of graphene: An introduction. Reviews of Modern Physics, 82 (3):2673–2700, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2673;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2673_1.

Perlmutter:2012:NLM

[Per12] Saul Perlmutter. Nobel Lecture: Measuring the acceleration of the cosmic expansion using supernovae. Reviews of Modern Physics, 84(3):1127–1149, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.84.
1127;   http://rmp.aps.org/abstract/RMP/v84/i3/p1127_
1;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/2011/.

Paladino:2014:NIS

[PGFA14] E. Paladino, Y. M. Galperin, G. Falci, and B. L. Altshuler. 1/f noise: Implications for solid-state quantum information. Reviews of Modern Physics, 86(2):361–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.86.361.

Presse:2013:PME

[PGLD13] Steve Pressé, Kingshuk Ghosh, Julian Lee, and Ken A. Dill. Principles of maximum entropy and maximum caliber in statistical physics. Reviews of Modern Physics, 85 (3):1115–1141, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1115;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1115_1.

Pradhan:2010:FPE

[PHC10] Srutarshi Pradhan, Alex Hansen, and Bikas K. Chakrabarti. Failure processes in elastic fiber bundles. Reviews of Modern Physics, 82(1):499–555, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.499;
http://rmp.aps.org/abstract/RMP/v82/i1/p499_1.

Phillips:2010:CIP

[Phi10] Philip Phillips. Colloquium: Identifying the propagating charge modes in doped Mott insulators. Reviews of Modern Physics, 82(2):1719–1742, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1719;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1719_1.

Pfutzner:2012:RDL

[PKGR12] M. Pfützner, M. Karny, L. V. Grigorenko, and K. Riisager. Radioactive decays at limits of nuclear stability. Reviews of Modern Physics, 84(2):567–619, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.567;
http://rmp.aps.org/abstract/RMP/v84/i2/p567_1.

Pellegrini:2016:PXR

[PMR16] C. Pellegrini, A. Marinelli, and S. Reiche. The physics of X-ray free-electron lasers. Reviews of Modern Physics, 88(1):015006– ??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.88.015006.

Pazourek:2015:ACP

[PNB15] Renate Pazourek, Stefan Nagele, and Joachim Burgdörfer. Attosecond chronoscopy of photoemission. Reviews of Modern Physics, 87(3):765–??, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.765.

Powers:2010:DFM

[Pow10a] Thomas R. Powers. Dynamics of filaments and membranes in a viscous fluid. Reviews of Modern Physics, 82(2):1607– 1631, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1607;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1607_1. See note

[Pow10b].

Powers:2010:PND

[Pow10b] Thomas R. Powers. Publisher’s note: Dynamics of filaments and membranes in a viscous fluid [Rev. Mod. Phys. 82, 1607 (2010)]. Reviews of Modern Physics, 82(2): 1945, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1945;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1945_1. See [Pow10a].

Poland:2019:CBT

[PRV19] David Poland, Slava Rychkov, and Alessandro Vichi. The conformal bootstrap: Theory, numerical techniques, and applications. Reviews of Modern Physics, 91(1):015002–??, January 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.91.015002.

Pastor-Satorras:2015:EPC

[PSCVV15] Romualdo Pastor-Satorras, Claudio Castellano, Piet Van Mieghem, and Alessandro Vespignani. Epidemic processes in complex networks. Reviews of Modern Physics, 87(3):925– ??, July/September 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.87.
925.

Pekola:2013:SEC

[PSM^{+}13] Jukka P. Pekola, Olli-Pentti Saira, Ville F. Maisi, Antti Kemppinen, Mikko Möttönen, Yuri A. Pashkin, and Dmitri V. Averin. Single-electron current sources: Toward a refined definition of the ampere. Reviews of Modern Physics, 85(4): 1421–1472, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1421;   http:
//rmp.aps.org/abstract/RMP/v85/i4/p1421_1.

Pezze:2018:QMN

[PSO^{+}18] Luca Pezzè, Augusto Smerzi, Markus K. Oberthaler, Roman Schmied, and Philipp Treutlein. Quantum metrology with nonclassical states of atomic ensembles. Reviews of Modern Physics, 90(3):035005–??, March 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.035005.

Polkovnikov:2011:CND

[PSSV11] Anatoli Polkovnikov, Krishnendu Sengupta, Alessandro Silva, and Mukund Vengalattore. Colloquium: Nonequilibrium dynamics of closed interacting quantum systems. Reviews of Modern Physics, 83(3):863–883, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.863;
http://rmp.aps.org/abstract/RMP/v83/i3/p863_1.

Parisi:2010:MFT

[PZ10] Giorgio Parisi and Francesco Zamponi. Mean-field theory of hard sphere glasses and jamming. Reviews of Modern Physics, 82(1):789–845, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.82.789;   http:
//rmp.aps.org/abstract/RMP/v82/i1/p789_1.

Qi:2011:TIS

[QZ11] Xiao-Liang Qi and Shou-Cheng Zhang. Topological insulators and superconductors. Reviews of Modern Physics, 83(4): 1057–1110, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1057;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1057_1.

Rowe:2012:DPS

[RCR12] D. J. Rowe, M. J. Carvalho, and J. Repka. Dual pairing of symmetry and dynamical groups in physics. Reviews of Modern Physics, 84(2):711–757, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.711;
http://rmp.aps.org/abstract/RMP/v84/i2/p711_1.

Ritsch:2013:CAC

[RDBE13] Helmut Ritsch, Peter Domokos, Ferdinand Brennecke, and Tilman Esslinger. Cold atoms in cavity-generated dynamical optical potentials. Reviews of Modern Physics, 85(2):553–601, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.553;   http://rmp.aps.org/
abstract/RMP/v85/i2/p553_1.

Rotter:2017:LFC

[RG17] Stefan Rotter and Sylvain Gigan. Light fields in complex media: Mesoscopic scattering meets wave control. Reviews of Modern Physics, 89(1):015005–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.015005.

Rohringer:2018:DRN

[RHT^{+}18] G. Rohringer, H. Hafermann, A. Toschi, A. A. Katanin, A. E. Antipov, M. I. Katsnelson, A. I. Lichtenstein, A. N. Rubtsov, and K. Held. Diagrammatic routes to nonlocal correlations beyond dynamical mean field theory. Reviews of Modern Physics, 90(2):025003–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.90.025003.

Riess:2012:NLM

[Rie12] Adam G. Riess. Nobel Lecture: My path to the accelerating Universe. Reviews of Modern Physics, 84(3):1165– 1175, July 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1165;   http:
//rmp.aps.org/abstract/RMP/v84/i3/p1165_1;   http://
www.nobelprize.org/nobel_prizes/physics/laureates/2011/

.

Rosenstein:2010:GLT

[RL10] Baruch Rosenstein and Dingping Li. Ginzburg–Landau theory of type II superconductors in magnetic field. Reviews of Modern Physics, 82(1):109–168, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.109;
http://rmp.aps.org/abstract/RMP/v82/i1/p109_1.

Ryd:2012:HDD

[RP12] Anders Ryd and Alexey A. Petrov. Hadronic D and D_{s} meson decays. Reviews of Modern Physics, 84(1):65–117, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.84.65;   http://rmp.aps.org/abstract/RMP/
v84/i1/p65_1.

Reiserer:2015:CBQ

[RR15] Andreas Reiserer and Gerhard Rempe. Cavity-based quantum networks with single atoms and optical photons. Reviews of Modern Physics, 87(4):1379–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.87.1379.

Rueff:2010:IXR

[RS10] Jean-Pascal Rueff and Abhay Shukla. Inelastic X-ray scattering by electronic excitations under high pressure. Reviews of Modern Physics, 82(1):847–896, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.82.847;
http://rmp.aps.org/abstract/RMP/v82/i1/p847_1.

Reineke:2013:WOL

[RTLL13] Sebastian Reineke, Michael Thomschke, Björn Lüssem, and Karl Leo. White organic light-emitting diodes: Status and perspective. Reviews of Modern Physics, 85(3):1245– 1293, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1245;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1245_1.

Rurali:2010:CSE

[Rur10] Riccardo Rurali. Colloquium: Structural, electronic, and transport properties of silicon nanowires. Reviews of Modern Physics, 82(1):427–449, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.82.427;   http:
//rmp.aps.org/abstract/RMP/v82/i1/p427_1.

Roy:2017:CSI

[RWF17] Dibyendu Roy, C. M. Wilson, and Ofer Firstenberg. Colloquium: Strongly interacting photons in one-dimensional continuum. Reviews of Modern Physics, 89(2):021001–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.021001.

Suter:2016:CPQ

[SÁ16] Dieter Suter and Gonzalo A. Álvarez. Colloquium: Protecting quantum information against environmental noise. Reviews of Modern Physics, 88(4):041001–??, October/December 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.041001.

Sagis:2011:DPI

[Sag11] Leonard M. C. Sagis. Dynamic properties of interfaces in soft matter: Experiments and theory. Reviews of Modern Physics, 83 (4):1367–1403, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1367;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1367_1.

Sarma:2011:ETT

[SAHR11] S. Das Sarma, Shaffique Adam, E. H. Hwang, and Enrico Rossi. Electronic transport in two-dimensional graphene. Reviews of Modern Physics, 83(2):407–470, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.407;
http://rmp.aps.org/abstract/RMP/v83/i2/p407_1.

Streltsov:2017:CQC

[SAP17] Alexander Streltsov, Gerardo Adesso, and Martin B. Plenio. Colloquium: Quantum coherence as a resource. Reviews of Modern Physics, 89(4):041003–??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.041003.

Safronova:2018:SNP

[SBD^{+}18] M. S. Safronova, D. Budker, D. DeMille, Derek F. Jackson Kimball, A. Derevianko, and Charles W. Clark. Search for new physics with atoms and molecules. Reviews of Modern Physics, 90(2):025008–??, February 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.025008.

Sato:2010:FPT

[SBK^{+}10] K. Sato, L. Bergqvist, J. Kudrnovský, P. H. Dederichs, O. Eriksson, I. Turek, B. Sanyal, G. Bouzerar, H. Katayama-Yoshida, V. A. Dinh, T. Fukushima, H. Kizaki, and R. Zeller. Firstprinciples theory of dilute magnetic semiconductors. Reviews of Modern Physics, 82(2):1633–1690, April 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.1633;   http://rmp.aps.org/abstract/RMP/v82/i2/p1633_
1.

Scalapino:2012:CTP

[Sca12] D. J. Scalapino. A common thread: The pairing interaction for unconventional superconductors. Reviews of Modern Physics, 84 (4):1383–1417, October 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.84.1383;   http:
//rmp.aps.org/abstract/RMP/v84/i4/p1383_1.

Schmidt:2012:NLA

[Sch12] Brian P. Schmidt. Nobel Lecture: Accelerating expansion of the universe through observations of distant supernovae. Reviews of Modern Physics, 84(3):1151–1163, July 2012. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
84.1151;   http://rmp.aps.org/abstract/RMP/v84/i3/p1151_
1;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/2011/.

Schellekens:2013:LIP

[Sch13] A. N. Schellekens. Life at the interface of particle physics and string theory. Reviews of Modern Physics, 85(4):1491– 1540, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1491;   http:
//rmp.aps.org/abstract/RMP/v85/i4/p1491_1.

Schliemann:2017:CPS

[Sch17] John Schliemann. Colloquium: Persistent spin textures in semiconductor nanostructures. Reviews of Modern Physics, 89(1): 011001–??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.89.
011001.

Shukla:2011:CNC

[SE11] P. K. Shukla and B. Eliasson. Colloquium: Nonlinear collective interactions in quantum plasmas with degenerate electron fluids. Reviews of Modern Physics, 83(3):885–906, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.83.885;   http://rmp.aps.org/
abstract/RMP/v83/i3/p885_1.

Sellwood:2014:SED

[Sel14] J. A. Sellwood. Secular evolution in disk galaxies. Reviews of Modern Physics, 86(1):1–??, January/March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.1.

Schardt:2010:HIT

[SESE10] Dieter Schardt, Thilo Elsässer, and Daniela Schulz-Ertner. Heavy-ion tumor therapy: Physical and radiobiological benefits. Reviews of Modern Physics, 82(1):383–425, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.383;   http://rmp.aps.org/
abstract/RMP/v82/i1/p383_1.

Sotiriou:2010:TG

[SF10a] Thomas P. Sotiriou and Valerio Faraoni. f(R) theories of gravity. Reviews of Modern Physics, 82(1):451–497, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.451;   http://rmp.aps.org/
abstract/RMP/v82/i1/p451_1.

Strandlie:2010:TVR

[SF10b] Are Strandlie and Rudolf Frühwirth. Track and vertex reconstruction: From classical to adaptive methods. Reviews of Modern Physics, 82(2):1419–1458, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1419;   http://rmp.aps.org/abstract/RMP/v82/i2/p1419_
1.

Sirlin:2013:RCP

[SF13] Alberto Sirlin and Andrea Ferroglia. Radiative corrections in precision electroweak physics: a historical perspective. Reviews of Modern Physics, 85(1):263–297, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.263;   http://rmp.aps.org/
abstract/RMP/v85/i1/p263_1.

Skryabin:2010:CLS

[SG10] Dmitry V. Skryabin and Andrey V. Gorbach. Colloquium: Looking at a soliton through the prism of optical supercontinuum. Reviews of Modern Physics, 82(2):1287– 1299, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1287;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1287_1.

Soriano:2013:CPD

[SGOMF13] Miguel C. Soriano, Jordi García-Ojalvo, Claudio R. Mirasso, and Ingo Fischer. Complex photonics: Dynamics and applications of delay-coupled semiconductors lasers. Reviews of Modern Physics, 85(1):421–470, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.85.421;
http://rmp.aps.org/abstract/RMP/v85/i1/p421_1.

Schumayer:2011:CPR

[SH11a] Dániel Schumayer and David A. W. Hutchinson. Colloquium: Physics of the Riemann hypothesis. Reviews of Modern Physics, 83(2):307–330, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.83.307;   http://rmp.
aps.org/abstract/RMP/v83/i2/p307_1. See note [SH11b].

Schumayer:2011:PNC

[SH11b] Dániel Schumayer and David A. W. Hutchinson. Publisher’s note: Colloquium: Physics of the Riemann hypothesis [Rev. Mod. Phys. 83, 307 (2011)]. Reviews of Modern Physics, 83(2): 769, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.83.769;   http://rmp.aps.
org/abstract/RMP/v83/i2/p769_1. See [SH11a].

Saito:2013:CHS

[SH13] Yukio Saito and Hiroyuki Hyuga. Colloquium: Homochirality: Symmetry breaking in systems driven far from equilibrium. Reviews of Modern Physics, 85(2):603–621, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.85.603;   http://rmp.aps.org/
abstract/RMP/v85/i2/p603_1.

Shuryak:2017:SCQ

[Shu17] Edward Shuryak. Strongly coupled quark-gluon plasma in heavy ion collisions. Reviews of Modern Physics, 89(3):035001–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.035001.

Spivak:2010:CTS

[SKKG10] B. Spivak, S. V. Kravchenko, S. A. Kivelson, and X. P. A. Gao. Colloquium: Transport in strongly correlated two dimensional electron fluids. Reviews of Modern Physics, 82 (2):1743–1766, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1743;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1743_1.

Stamper-Kurn:2013:SBG

[SKU13] Dan M. Stamper-Kurn and Masahito Ueda. Spinor Bose gases: Symmetries, magnetism, and quantum dynamics. Reviews of Modern Physics, 85(3):1191–1244, July 2013. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
85.1191;   http://rmp.aps.org/abstract/RMP/v85/i3/p1191_
1.

Smith:2010:NLI

[Smi10] George E. Smith. Nobel Lecture: The invention and early history of the CCD. Reviews of Modern Physics, 82(3): 2307–2312, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2307;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2307_1.

Ma:2016:DCG

[sMKZ16] Xiao song Ma, Johannes Kofler, and Anton Zeilinger. Delayedchoice gedanken experiments and their realizations. Reviews of Modern Physics, 88(1):015005–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.88.015005.

Singer:2010:CTI

[SPM^{+}10] Kilian Singer, Ulrich Poschinger, Michael Murphy, Peter Ivanov, Frank Ziesel, Tommaso Calarco, and Ferdinand Schmidt-Kaler. Colloquium: Trapped ions as quantum bits: Essential numerical tools. Reviews of Modern Physics, 82(3): 2609–2632, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2609;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2609_1.

Stankovski:2017:CFU

[SPMS17] Tomislav Stankovski, Tiago Pereira, Peter V. E. McClintock, and Aneta Stefanovska. Coupling functions: Universal insights into dynamical interaction mechanisms. Reviews of Modern Physics, 89(4):045001–??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.89.045001.

Sprouse:2014:EPT

[Spr14] Gene D. Sprouse. Editorial: PRX takes on a new role. Reviews of Modern Physics, 86(4):1187–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.1187.

Saarikoski:2010:VQD

[SRHM10] H. Saarikoski, S. M. Reimann, A. Harju, and M. Manninen. Vortices in quantum droplets: Analogies between boson and fermion systems. Reviews of Modern Physics, 82 (3):2785–2834, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2785;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2785_1.

Schwarz:2013:PAC

[SS13] Ulrich S. Schwarz and Samuel A. Safran. Physics of adherent cells. Reviews of Modern Physics, 85(3):1327– 1381, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1327;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1327_1.

Sangouard:2011:QRB

[SSdRG11] Nicolas Sangouard, Christoph Simon, Hugues de Riedmatten, and Nicolas Gisin. Quantum repeaters based on atomic ensembles and linear optics. Reviews of Modern Physics, 83 (1):33–80, January 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.83.33;   http:/
/rmp.aps.org/abstract/RMP/v83/i1/p33_1.

Seiden:2011:CSP

[ST11] G. Seiden and P. J. Thomas. Complexity, segregation, and pattern formation in rotating-drum flows. Reviews of Modern Physics, 83(4):1323–1365, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.83.
1323;   http://rmp.aps.org/abstract/RMP/v83/i4/p1323_
1.

Salter:2018:ETT

[ST18] Matthew Salter and Michael Thoennessen. Editorial: A tale of two anniversaries: 125 years of the Physical Review and 25 years of Physical Review E. Reviews of Modern Physics, 90(1):010001–??, January 2018. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.90.010001.

Stewart:2011:SIC

[Ste11] G. R. Stewart. Superconductivity in iron compounds. Reviews of Modern Physics, 83(4):1589–1652, October 2011. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
83.1589;   http://rmp.aps.org/abstract/RMP/v83/i4/p1589_
1.

Strickland:2019:NLG

[Str19] Donna Strickland. Nobel Lecture: Generating high-intensity ultrashort optical pulses. Reviews of Modern Physics, 91(3):

030502–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.91.
030502.

Sun:2014:IFD

[Sun14] Mao Sun. Insect flight dynamics: Stability and control. Reviews of Modern Physics, 86(2):615–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.86.615.

Sinova:2015:SHE

[SVW^{+}15] Jairo Sinova, Sergio O. Valenzuela, J. Wunderlich, C. H. Back, and T. Jungwirth. Spin Hall effects. Reviews of Modern Physics, 87(4):1213–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1213.

Snijders:2010:CEI

[SW10] Paul C. Snijders and Hanno H. Weitering. Colloquium: Electronic instabilities in self-assembled atom wires. Reviews of Modern Physics, 82(1):307–329, January 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
82.307;   http://rmp.aps.org/abstract/RMP/v82/i1/p307_
1.

Sefkow:2016:ETP

[SWK^{+}16] Felix Sefkow, Andy White, Kiyotomo Kawagoe, Roman Pöschl, and José Repond. Experimental tests of particle flow calorimetry. Reviews of Modern Physics, 88(1):015003–??, January 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.015003.

Saffman:2010:QIR

[SWM10] M. Saffman, T. G. Walker, and K. Mølmer. Quantum information with Rydberg atoms. Reviews of Modern Physics, 82 (3):2313–2363, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.2313;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p2313_1.

Sahu:2019:CIP

[SZ19] Subin Sahu and Michael Zwolak. Colloquium: Ionic phenomena in nanoscale pores through 2D materials. Reviews of Modern Physics, 91(2):021004–??, February 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.021004.

Tang:2018:CCD

[TB18] Evelyn Tang and Danielle S. Bassett. Colloquium: Control of dynamics in brain networks. Reviews of Modern Physics, 90 (3):031003–??, March 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.90.
031003.

Tanaka:2016:SPW

[TBC^{+}16] H. Tanaka, M. J. Brunger, L. Campbell, H. Kato, M. Hoshino, and A. R. P. Rau. Scaled plane-wave Born cross sections for atoms and molecules. Reviews of Modern Physics, 88(2):025004– ??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.88.025004.

Terhal:2015:QEC

[Ter15] Barbara M. Terhal. Quantum error correction for quantum memories. Reviews of Modern Physics, 87(2):307–??, April/ June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.87.307.

Takezoe:2010:ALC

[TGC10] Hideo Takezoe, Ewa Gorecka, and Mojca Cepic. Antiferroelectric liquid crystals: Interplay of simplicity and complexity. Reviews of Modern Physics, 82(1):897–937, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.897;   http://rmp.aps.org/
abstract/RMP/v82/i1/p897_1.

Thorne:2018:NLL

[Tho18] Kip S. Thorne. Nobel lecture: LIGO and gravitational waves III. Reviews of Modern Physics, 90(4):040503–??, April 2018.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL https://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.040503.

Tohsaki:2017:CSP

[THSR17] Akihiro Tohsaki, Hisashi Horiuchi, Peter Schuck, and Gerd Röpke. Colloquium: Status of α-particle condensate structure of the Hoyle state. Reviews of Modern Physics, 89(1):011002– ??, January 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.w3.org/

1998/Math/MathML.

Tielens:2013:MU

[Tie13] A. G. G. M. Tielens. The molecular universe. Reviews of Modern Physics, 85(3):1021–1081, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.85.
1021;   http://rmp.aps.org/abstract/RMP/v85/i3/p1021_
1.

Tomza:2019:CHI

[TJG^{+}19] Michał Tomza, Krzysztof Jachymski, Rene Gerritsma, Antonio Negretti, Tommaso Calarco, Zbigniew Idziaszek, and Paul S. Julienne. Cold hybrid ion–atom systems. Reviews of Modern Physics, 91(3):035001–??, March 2019. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.91.035001.

Tuomisto:2013:DIS

[TM13] Filip Tuomisto and Ilja Makkonen. Defect identification in semiconductors with positron annihilation: Experiment and theory. Reviews of Modern Physics, 85(4):1583–1631, October 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.85.1583;   http://rmp.aps.org/abstract/RMP/
v85/i4/p1583_1.

Torquato:2010:JHP

[TS10a] S. Torquato and F. H. Stillinger. Jammed hard-particle packings: From Kepler to Bernal and beyond. Reviews of Modern Physics, 82(3):2633–2672, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
2633;   http://rmp.aps.org/abstract/RMP/v82/i3/p2633_

1. See note [TS10b].

Torquato:2010:PNJ

[TS10b] S. Torquato and F. H. Stillinger. Publisher’s note: Jammed hard-particle packings: From Kepler to Bernal and beyond [Rev. Mod. Phys. 82, 2633 (2010)]. Reviews of Modern Physics, 82(4):3197, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.3197;   http:
//rmp.aps.org/abstract/RMP/v82/i4/p3197_1. See [TS10a].

Takabatake:2014:PGE

[TSNK14a] Toshiro Takabatake, Koichiro Suekuni, Tsuneyoshi Nakayama, and Eiji Kaneshita. Phonon-glass electron-crystal thermoelectric clathrates: Experiments and theory. Reviews of Modern Physics, 86(2):669–??, April/June 2014. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.86.669. See note [TSNK14b].

Takabatake:2014:PNP

[TSNK14b] Toshiro Takabatake, Koichiro Suekuni, Tsuneyoshi Nakayama, and Eiji Kaneshita. Publisher’s note: Phonon-glass electroncrystal thermoelectric clathrates: Experiments and theory [Rev. Mod. Phys. 86, 669 (2014)]. Reviews of Modern Physics, 86 (2):841–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.86.

841. See [TSNK14a].

Turner:2010:VCS

[TVN10] Ari M. Turner, Vincenzo Vitelli, and David R. Nelson. Vortices on curved surfaces. Reviews of Modern Physics, 82 (2):1301–1348, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1301;   http:
//rmp.aps.org/abstract/RMP/v82/i2/p1301_1.

Ubachs:2016:CSD

[UBS^{+}16] W. Ubachs, J. Bagdonaite, E. J. Salumbides, M. T. Murphy, and L. Kaper. Colloquium: Search for a drifting proton–electron mass ratio from H_{2}. Reviews of Modern Physics, 88(2):021003– ??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.w3.org/

1998/Math/MathML.

Ulbricht:2011:CDS

[UHS^{+}11] Ronald Ulbricht, Euan Hendry, Jie Shan, Tony F. Heinz, and Mischa Bonn. Carrier dynamics in semiconductors studied with time-resolved terahertz spectroscopy. Reviews of Modern Physics, 83(2):543–586, April 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.543;
http://rmp.aps.org/abstract/RMP/v83/i2/p543_1. See er-

ratum [UHS^{+}17].

Ulbricht:2017:ECD

[UHS^{+}17] Ronald Ulbricht, Euan Hendry, Jie Shan, Tony F. Heinz, and Mischa Bonn. Erratum: Carrier dynamics in semiconductors studied with time-resolved terahertz spectroscopy [Rev. Mod. Phys. 83, 543 (2011)]. Reviews of Modern Physics, 89(2):029901–??, February 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.89.029901. See [UHS^{+}11].

Urbaszek:2013:NSP

[UMA^{+}13] Bernhard Urbaszek, Xavier Marie, Thierry Amand, Olivier Krebs, Paul Voisin, Patrick Maletinsky, Alexander Högele, and Atac Imamoglu. Nuclear spin physics in quantum dots: An optical investigation. Reviews of Modern Physics, 85 (1):79–133, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.85.79;   http:/
/rmp.aps.org/abstract/RMP/v85/i1/p79_1.

Varoquaux:2015:ACF

[Var15] Eric Varoquaux. Anderson’s considerations on the flow of superfluid helium: Some offshoots. Reviews of Modern Physics, 87(3):803–??, July/September 2015. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.87.803.

Vassen:2012:CTM

[VCTL^{+}12] Wim Vassen, Claude Cohen-Tannoudji, Michele Leduc, Denis Boiron, Christoph I. Westbrook, Andrew Truscott, Ken Baldwin, Gerhard Birkl, Pablo Cancio, and Marek Trippenbach. Cold and trapped metastable noble gases. Reviews of Modern Physics, 84(1):175–210, January 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.175;
http://rmp.aps.org/abstract/RMP/v84/i1/p175_1.

Varlamov:2018:FSR

[VGG18] A. A. Varlamov, A. Galda, and A. Glatz. Fluctuation spectroscopy: From Rayleigh–Jeans waves to Abrikosov vortex clusters. Reviews of Modern Physics, 90(1):015009–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.90.015009.

vanHouselt:2010:CTR

[vHZ10] Arie van Houselt and Harold J. W. Zandvliet. Colloquium: Time-resolved scanning tunneling microscopy. Reviews of Modern Physics, 82(2):1593–1605, April 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
1593;   http://rmp.aps.org/abstract/RMP/v82/i2/p1593_
1.

Vanossi:2013:CMF

[VMU^{+}13] Andrea Vanossi, Nicola Manini, Michael Urbakh, Stefano Zapperi, and Erio Tosatti. Colloquium: Modeling friction: From nanoscale to mesoscale. Reviews of Modern Physics, 85(2):529– 552, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.85.529;   http://rmp.aps.
org/abstract/RMP/v85/i2/p529_1.

Vitanov:2017:SRA

[VRSB17] Nikolay V. Vitanov, Andon A. Rangelov, Bruce W. Shore, and Klaas Bergmann. Stimulated Raman adiabatic passage in physics, chemistry, and beyond. Reviews of Modern Physics, 89(1):015006–??, January 2017. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.89.015006.

vonToussaint:2011:BIP

[vT11] Udo von Toussaint. Bayesian inference in physics. Reviews of Modern Physics, 83(3):943–999, July 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.83.943;
http://rmp.aps.org/abstract/RMP/v83/i3/p943_1.

Vos:2015:SVN

[VWT15] K. K. Vos, H. W. Wilschut, and R. G. E. Timmermans. Symmetry violations in nuclear and neutron β decay. Reviews of Modern Physics, 87(4):1483–??, October/December 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.1483.

Watts:2016:CMN

[WAC^{+}16] Anna L. Watts, Nils Andersson, Deepto Chakrabarty, Marco Feroci, Kai Hebeler, Gianluca Israel, Frederick K. Lamb, M. Coleman Miller, Sharon Morsink, Feryal Özel, Alessandro Patruno, Juri Poutanen, Dimitrios Psaltis, Achim Schwenk, Andrew W. Steiner, Luigi Stella, Laura Tolos, and Michiel van der Klis. Colloquium: Measuring the neutron star equation of state using Xray timing. Reviews of Modern Physics, 88(2):021001–??, February 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.88.021001.

Wang:2018:CEA

[WCG^{+}18] Gang Wang, Alexey Chernikov, Mikhail M. Glazov, Tony F. Heinz, Xavier Marie, Thierry Amand, and Bernhard Urbaszek. Colloquium: Excitons in atomically thin transition metal dichalcogenides. Reviews of Modern Physics, 90(2):021001–??, February 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.021001.

Woods:2016:MPC

[WDT^{+}16] L. M. Woods, D. A. R. Dalvit, A. Tkatchenko, P. Rodriguez-Lopez, A. W. Rodriguez, and R. Podgornik. Materials perspective on Casimir and van der Waals interactions. Reviews of Modern Physics, 88(4):045003–??, October/December 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.045003.

Weiss:2018:NLL

[Wei18] Rainer Weiss. Nobel lecture: LIGO and the discovery of gravitational waves I. Reviews of Modern Physics, 90(4):040501– ??, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL https://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.90.040501.

Wen:2017:CZQ

[Wen17] Xiao-Gang Wen. Colloquium: Zoo of quantum-topological phases of matter. Reviews of Modern Physics, 89(4):041004– ??, October/December 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.89.
041004.

West:2014:CFC

[Wes14] Bruce J. West. Colloquium: Fractional calculus view of complexity: A tutorial. Reviews of Modern Physics, 86(4):1169–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.1169.

Wietfeldt:2011:CNL

[WG11] Fred E. Wietfeldt and Geoffrey L. Greene. Colloquium: The neutron lifetime. Reviews of Modern Physics, 83(4):1173– 1192, October 2011. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.83.1173;   http:
//rmp.aps.org/abstract/RMP/v83/i4/p1173_1.

Wineland:2013:NLS

[Win13] David J. Wineland. Nobel Lecture: Superposition, entanglement, and raising Schrödinger’s cat. Reviews of Modern Physics, 85(3):1103–1114, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.85.1103;   http:
//rmp.aps.org/abstract/RMP/v85/i3/p1103_1.

Witten:2016:FPI

[Wit16] Edward Witten. Fermion path integrals and topological phases. Reviews of Modern Physics, 88(3):035001–??, July/September 2016. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.88.035001.

Witten:2018:AME

[Wit18] Edward Witten. APS Medal for Exceptional Achievement in Research: Invited article on entanglement properties of quantum field theory. Reviews of Modern Physics, 90(4):045003:1– 045003:38, April 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

journals.aps.org/rmp/abstract/10.1103/RevModPhys.90.
045003.

Weedbrook:2012:GQI

[WPGP^{+}12] Christian Weedbrook, Stefano Pirandola, Raúl García-Patrón, Nicolas J. Cerf, Timothy C. Ralph, Jeffrey H. Shapiro, and Seth Lloyd. Gaussian quantum information. Reviews of Modern Physics, 84(2):621–669, April 2012. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.84.621;
http://rmp.aps.org/abstract/RMP/v84/i2/p621_1.

Xiang:2013:HQC

[XAYN13] Ze-Liang Xiang, Sahel Ashhab, J. Q. You, and Franco Nori. Hybrid quantum circuits: Superconducting circuits interacting with other quantum systems. Reviews of Modern Physics, 85 (2):623–653, April 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.85.623;   http://rmp.
aps.org/abstract/RMP/v85/i2/p623_1.

Xiao:2010:BPE

[XCN10] Di Xiao, Ming-Che Chang, and Qian Niu. Berry phase effects on electronic properties. Reviews of Modern Physics, 82 (3):1959–2007, July 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.82.1959;   http:
//rmp.aps.org/abstract/RMP/v82/i3/p1959_1.

Yamada:2010:MR

[YKJ10] Masaaki Yamada, Russell Kulsrud, and Hantao Ji. Magnetic reconnection. Reviews of Modern Physics, 82(1):603–664, January 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.82.603;   http://rmp.aps.org/
abstract/RMP/v82/i1/p603_1.

Zhitomirsky:2013:CSM

[ZC13] M. E. Zhitomirsky and A. L. Chernyshev. Colloquium: Spontaneous magnon decays. Reviews of Modern Physics, 85(1): 219–242, January 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.85.219;   http://rmp.
aps.org/abstract/RMP/v85/i1/p219_1.

Zaburdaev:2015:LW

[ZDK15] V. Zaburdaev, S. Denisov, and J. Klafter. Lévy walks. Reviews of Modern Physics, 87(2):483–??, April/June 2015. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.87.483.

Zwanenburg:2013:SQE

[ZDM^{+}13] Floris A. Zwanenburg, Andrew S. Dzurak, Andrea Morello, Michelle Y. Simmons, Lloyd C. L. Hollenberg, Gerhard Klimeck, Sven Rogge, Susan N. Coppersmith, and Mark A. Eriksson. Silicon quantum electronics. Reviews of Modern Physics, 85(3): 961–1019, July 2013. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.85.961;   http://rmp.
aps.org/abstract/RMP/v85/i3/p961_1.

Zapf:2014:BEC

[ZJB14a] Vivien Zapf, Marcelo Jaime, and C. D. Batista. Bose– Einstein condensation in quantum magnets. Reviews of Modern Physics, 86(2):563–??, April/June 2014. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.86.563. See erratum [ZJB14b].

Zapf:2014:EBE

[ZJB14b] Vivien Zapf, Marcelo Jaime, and C. D. Batista. Erratum: Bose– Einstein condensation in quantum magnets [Rev. Mod. Phys. 86, 563 (2014)]. Reviews of Modern Physics, 86(4):1453–??, April/June 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.86.1453. See

[ZJB14a].

Zhou:2017:QSL

[ZKN17] Yi Zhou, Kazushi Kanoda, and Tai-Kai Ng. Quantum spin liquid states. Reviews of Modern Physics, 89(2):025003–??, February 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.89.025003.

Zeravcic:2017:CTL

[ZMB17] Zorana Zeravcic, Vinothan N. Manoharan, and Michael P. Brenner. Colloquium: Toward living matter with colloidal particles. Reviews of Modern Physics, 89(3):031001–??, March 2017. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.89.031001.